### v3

# Kac–Moody Symmetric Spaces: arbitary symmetrizable complex or almost split real type

Abstract
We revisit the theory of Kac–Moody symmetric spaces in order to include the affine case, complex numbers, and Galois descent to almost split real Kac–Moody symmetric spaces.
1 Introduction
Kac–Moody groups have been studied extensively since the 1980s, initiated by [KP85] and [Tit87]. However, for some reason Kac–Moody symmetric spaces have only been formally introduced in [FHHK20], although the heuristic idea that Kac–Moody symmetric spaces might be useful has been around in particular in mathematical physics for some time.
The focus of [FHHK20] was on symmetric spaces for split real Kac–Moody groups modelled with maximal flats of minimal possible dimension in order to propose geometries that might allow a primitive action as in [Cap05] – as this might then help prove abstract simplicity of the acting group. It is thus not surprising that the methods of [FHHK20] failed for affine Kac–Moody groups.
In the present article we vary on the theme of Kac–Moody symmetric spaces and introduce models with higher-dimensional flats that finally provide examples of Kac–Moody symmetric spaces also in the affine case. The key difference in the modelling can actually be found already on Kac–Moody algebra level and is described in Remark 3.3 and Convention 3.4. Not surprising, it hence turns out that the Kac–Moody symmetric spaces introduced here are identical to the ones introduced in [FHHK20] if and only if the generalized Cartan matrix is invertible (see Section 7).
Moreover, in the present article we also include the field $\mathbb{C}$ of complex numbers (see Convention 3.5), and we conduct Galois descent based on [Rém02], in order to model Kac–Moody symmetric spaces also in the almost split real case; see Section 9.
We took quite some effort to make the present article accessible independent from [FHHK20], but it is quite natural that we need to borrow and cite a wealth of results from [FHHK20]; on top of that, also a detailed comparison of our results with the results from [FHHK20] is called for.
We stress again that the purpose of this article is to provide an alternative definition in order to approach Kac–Moody symmetric spaces, that allows to also cover the affine case. On top of that we deal with the complex case and the almost split real case.
Our main result is the following:
Theorem 9.2
Let $\mathbb{K}\in\{\mathbb{R},\mathbb{C}\}$ and let $G$ be an (almost) split Kac–Moody $\mathbb{K}$-group satisfying (8.2) and denote by $G(\mathbb{K})$ the rational points of $G$. The quotient
$X(\mathbb{K})\coloneqq\nicefrac{{G(\mathbb{K})}}{{K(\mathbb{K})}}$
is a symmetric space in the sense of Loos [Loo69] with the reflection map $\mu^{\operatorname{as}}$. Moreover, there is a natural action of $G(\mathbb{K})$ on the symmetric space given by automorphisms
$G(\mathbb{K})$ $\to\operatorname{Sym}(X(\mathbb{K}))$
$g$ $\mapsto(hK(\mathbb{K})\mapsto ghK(\mathbb{K})).$
Acknowledgment:
The authors thank the DFG SPP 2026 priority program “Geometry at infinity” for partial financial support via KO 4323/14. The first half of the project, both authors were affiliated with and funded by JLU Giessen, the second half affiliated with and funded by CAU Kiel. The authors would also like to thank Paul Zellhofer in particular for a wealth of very helpful discussions and comments. This article is based on the PhD thesis of the second author that has been advised by the first author.
2 Basics on Symmetric Spaces
In [Loo69], Loos analyzes Riemannian and pseudo-Riemannian symmetric spaces and states an abstract definition without any use of differential geometry. This makes the definition amenable to situations in which one tries to generalize symmetric spaces into a context without a differential structure.
Definition 2.1 (Definition 1, Chapter II in [Loo69])
A topological space $X$ with a continuous map, which is called reflection, $\mu\colon X\times X\to X$, $(x,y)\mapsto\mu(x,y)\coloneqq x.y$ is called a symmetric space $(X,\mu)$ if the map satisfies the following properties:
  1. S.1
    $x.x=x$ for all $x\in X$,
  2. S.2
    $x.(x.y)=y$ for all $x,y\in X$,
  3. S.3
    $x.(y.z)=(x.y).(x.z)$ for all $x,y,z\in X$, and
  4. S.4
    for every $x\in X$ there is a neighborhood $U$ such that $x.y=y$ implies $x=y$ for all $y\in U$.
If the Riemannian symmetric space is of non-compact type, i.e. $X$ is non-compact, one can use the global condition instead of the local condition S.4:
$\begin{matrix}\textbf{S.4}_{\text{global}}&\forall x,y\in X\;\;y=y\Longrightarrow x=y.\end{matrix}$
[…]
A morphism of symmetric spaces is a (continuous) map $f\colon X_{1}\to X_{2}$ between two symmetric spaces $(X_{1},\mu_{1})$ and $(X_{2},\mu_{2})$ satisfying $f(\mu_{1}(x,y))=\mu_{2}(f(x),f(y))$. A pair $((X,\mu),\beta)$ consisting of a symmetric space $(X,\mu)$ and a point $\beta\in X$ is called a pointed symmetric space. A morphism between pointed symmetric spaces is a morphism between their underlying symmetric spaces with the additional property that the base point is preserved.
In the rest of the article, if the map $\mu$ is clear from the context, we will write $X$ for the symmetric space pair $(X,\mu)$ for short.
Example 2.2
  1. 1
    One of the first examples that comes to mind of a symmetric space is the $n$-dimensional Euclidean space $\mathbb{E}^{n}$ for any $n\in\mathbb{N}$. In particular, the symmetric space is given by the pair $\mathbb{R}^{n}$ together with the point reflection on $x\in\mathbb{R}^{n}$, i.e. the symmetric space map $\mu_{\text{Eucl}}$ is defined as $\mu_{\text{Eucl}}(x,y)\coloneqq 2x-y$.
  2. 2
    Considering the sphere with the corresponding geodesic reflection leads to an example, which satisfies only S.4.
  3. 3
[…]
  4. 4
    Another example is any topological group $G$ with the reflection $\mu\colon G\times G\to G$, $\mu(g,h)\coloneqq hgh^{-1}$. Here S.4 is in general not satisfied, but one can easily check that S.1 - S.3 are satisfied. For a detailed calculation of the first three axioms for topological groups, we refer to the proof of [FHHK20, Porposition 4.2]. There, an assumption is made that guarantees the fourth axiom, and this assumption also holds for the Kac–Moody situation, as we will see in 4.3.
Define the point reflection at $x\in X$ as the map
$s_{x}\colon X\to X\,,\;y\mapsto\mu(x,y).$
The product of two point reflections is called transvection and the group
$\operatorname{Trans}(X)\coloneqq\langle s_{x}\circ s_{y}\mid x,y\in X\rangle$
is called the transvection group of $X$.
Lemma 2.3 (p.64, line 15 of [Loo69])
Let $(X,\mu)$ be a symmetric space, $x$ a point in $X$ and let $\alpha\in\operatorname{Aut}(X)$ be an automorphism of the symmetric space. Then
$\alpha\circ s_{x}\circ\alpha^{-1}=s_{\alpha(x)}.$
Proof:
For each $p\in X$ one calculates
$\left(\alpha\circ s_{x}\circ\alpha^{-1}\right)(p)=\alpha\left(s_{x}\left(\alpha^{-1}(p)\right)\right)=\alpha\left(\mu\left(x,\alpha^{-1}(p)\right)\right)=\mu(\alpha(x),p)=s_{\alpha(x)}(p).$
By setting $\alpha=s_{y}$ for $y\in X$ in the previous lemma, we obtain the useful identity
$s_{y}\circ s_{x}\circ s_{y}^{-1}=s_{y}\circ s_{x}\circ s_{y}=s_{s_{y}(x)}.$ (1)
Remark 2.4
Consider a pointed symmetric space $(X,\mu)$ with base point $b\in X$. We now want to describe the symmetric space via the quadratic representation, see [Loo69, Section II.1] or [FHHK20, Remark 2.10]. For $x\in X$ define
$t_{x}\coloneqq s_{x}\circ s_{b}\in\operatorname{Trans}(X)$
and
$T(X,b)\coloneqq\{t_{x}\mid x\in X\}.$
Evidently, the map
$\eta\colon X\to T(X,b)\,,\;x\mapsto t_{x}$
is surjective. Since $t_{x}\circ s_{b}=s_{x}$ and since $s_{x}=s_{y}$ implies $x=y$, the map $\eta$ is also injective.
Model a symmetric space on the set $T(X,b)$ by defining the reflection map
$\mu_{T}\colon T(X,b)\times T(X,b)\to T(X,b)\,,\;(s,t)\mapsto s\circ t^{-1}\circ s.$
A direct calculation shows that $\eta:(X,\mu)\to(T(X,b),\mu_{T})$ is an isomorphism of symmetric spaces. Therefore, one has to check that the reflection map of $(X,\mu)$ is preserved, i.e. $\eta(\mu(x,y))=\mu_{T}(t_{x},t_{y})$. On the right-hand side one can calculate
$\mu_{T}(t_{x},t_{y})$ $=t_{x}\circ t_{y}^{-1}\circ t_{x}$
$=(s_{x}\circ s_{b})\circ(s_{y}\circ s_{b})^{-1}\circ(s_{x}\circ s_{b})$
$=s_{x}\circ s_{b}\circ s_{b}\circ s_{y}\circ s_{x}\circ s_{b}$
$=s_{x}\circ s_{y}\circ s_{x}\circ s_{b}$
[…]
And on the left-hand side one has
$\eta(\mu(x,y))=\eta(s_{x}(y))=t_{s_{x}(y)}.$
Recall from [FHHK20, Section 2.15] the following useful definitions and properties of an abstract symmetric space $(X,\mu)$:
  • A midpoint of two points $x,y\in X$ is a point $m\in X$, such that $\mu(m,x)=y$ and $\mu(m,y)=x$.
  • Call a subspace $U\subseteq X$
    • a reflection subspace, if for all pair of points $x,y\in U$ the image $\mu(x,y)$ is contained in $U$.
    • midpoint convex, if for all $x,y\in U$ there is a midpoint of $x$ and $y$ in $U$.
  • Two points $x,y\in X$ weakly commute, if for every point $p\in X$ applies $x.(p.(y.p))=y.(p.(x.p))$.
  • Two points $x,y\in X$ commute, if for all points $p,q\in X$ applies $x.(p.(y.q))=y.(p.(x.q))$.
  • A reflection subspace $F\subseteq X$ is a (weak) flat, if it is closed, midpoint convex, contains at least two points and if all points $x,y\in F$ (weakly) commute.
Definition 2.5 (cp. Definition 2.23 from [FHHK20])
A Euclidean flat of rank $n$ is a closed symmetric subspace $F\subseteq X$ isomorphic to the $n$-dimensional Euclidean space $\mathbb{E}^{n}$ as a symmetric space. For $n=1$, the flat is called a geodesic.
3 Kac–Moody Groups
In this section, we give a brief overview of the basic notions as well as the construction of a Kac–Moody group. Most of this can be found in [Kac90], [Tit87], [Rém02] or [Mar18].
3.1 Kac–Moody Algebra
Definition 3.1
Let $\mathbb{A}=\left(a_{ij}\right)_{1\leq i,j\leq n}\in\mathbb{Z}^{n\times n}$ be a square matrix over the integers. The matrix $\mathbb{A}$ is called a generalized Cartan matrix if it satisfies the following properties
  • $a_{ii}=2$ for all $i=1,\ldots,n$,
  • $a_{ij}\leq 0$ for all $i,j=1,\ldots,n$, $i\neq j$, and
  • $a_{ij}=0$ if and only if $a_{ji}=0$.
The generalized Cartan matrix $\mathbb{A}$ is indecomposable if it cannot be decomposed into a non-trivial direct sum after rearranging the indices, cf. [Kac90, §1.1]. Furthermore, $\mathbb{A}$ is symmetrizable if there is a symmetric matrix $B\in\mathbb{R}^{n\times n}$ and an invertible diagonal matrix $D\in(\mathbb{R}^{\times})^{n\times n}$ such that $\mathbb{A}=DB$, cf. [Kac90, §2.1].
If a generalized Cartan matrix is the Cartan matrix of a finite-dimensional semisimple complex Lie algebra, then $\mathbb{A}$ is of spherical type; otherwise, $\mathbb{A}$ is non-spherical.
In [Kac90] the notion of a realization of $\mathbb{A}$ is introduced, which assigns a triple to a generalized Cartan matrix, allowing the construction of a Kac–Moody algebra. As in the finite-dimensional situation, a Kac–Moody algebra is naturally related to several different groups; therefore, in order to define a Kac–Moody group associated to the algebra one needs a more subtle definition (see [Rém02, Chapitre 8] or [Mar18, Definition 7.9]).
Definition 3.2
Let $\mathbb{A}=\left(a_{ij}\right)_{i,j\in I}$ be a generalized Cartan matrix with index set $I$, let $\Lambda$ be a free $\mathbb{Z}$-module of finite rank and $\Lambda^{\vee}$ the $\mathbb{Z}$-dual of $\Lambda$, and let $(c_{i})_{i\in I}\subset\Lambda$ and $(h_{i})_{i\in I}\subset\Lambda^{\vee}$ such that $c_{j}(h_{i})=a_{ij}$. The quintuple $\mathcal{D}\coloneqq\left(I,\mathbb{A},\Lambda,(c_{i})_{i\in I},(h_{i})_{i\in I}\right)$ is called a Kac–Moody root datum.
Consider a generalized Cartan matrix $\mathbb{A}$ of arbitrary type and a corresponding Kac–Moody root datum $\mathcal{D}$. Define the Cartan subalgebra
$\mathfrak{h}_{\mathcal{D}}\coloneqq\Lambda^{\vee}\otimes_{\mathbb{Z}}\mathbb{C}$
with elements, $\alpha_{i}^{\vee}\coloneqq h_{i}\otimes 1$ and define the dual space
$\mathfrak{h}_{\mathcal{D}}^{*}\coloneqq\Lambda\otimes_{\mathbb{Z}}\mathbb{C}$
with elements $\alpha_{i}\coloneqq c_{i}\otimes 1$. Call the elements $\Pi=\{\alpha_{i}\mid i\in I\}\subset\mathfrak{h}^{*}$ simple roots, $\Pi^{\vee}=\{\alpha_{i}^{\vee}\mid i\in I\}\subset\mathfrak{h}$ simple coroots and denote with $\Delta^{re}$ the set of real roots (cf. [Kac90, §5.1] or [Mar18, Chapter 6.1]).
Remark 3.3
According to the approach proposed in Kac’s book [Kac90, §1.1], one wants the sets $\Pi$ and $\Pi^{\vee}$ to be linearly independent. Therefore, in the case that $\mathbb{A}$ is of affine type it follows from the relation $\langle\alpha_{i},\alpha_{j}^{\vee}\rangle=a_{ij}\in\mathbb{A}$, that $\Pi$ is linearly dependent. Hence, extending the Cartan subalgebra $\mathfrak{h}$ solves this problem, as done in [Kac90, Proof of Proposition 1.1] and [Mar18, Chapter 3.5 and Example 7.10]: Let $\mathbb{A}\in\mathbb{Z}^{n\times n}$ and $l$ be the rank of $\mathbb{A}$, then the extended Cartan subalgebra must be $\mathfrak{h}\cong\mathbb{C}^{2n-l}$. The dual space is also extended and constructed so that the roots still satisfy the relation to $\mathbb{A}$.
Let us now denote by $\mathcal{D}_{\mathrm{Kac}}^{\mathbb{A}}$ the root datum where the elements $c_{i}$ and $h_{i}$ are linearly independent and denote by $\Lambda_{\mathrm{Kac}}$ the corresponding $\mathbb{Z}$-module, which has minimum rank with respect to these properties, i.e. the rank of $\Lambda_{\mathrm{Kac}}$ is equal to $2n-l$. Similarly, we use $\Lambda_{\mathrm{Kac}}^{\vee}$ to denote the $\mathbb{Z}$-dual. Moreover, denote by $\{v_{i}\mid 1\leq i\leq 2n-l\}$ the $\mathbb{Z}$-basis of the extended $\Lambda_{\mathrm{Kac}}$, where $h_{i}=v_{i}$ for $1\leq i\leq n$. We follow here the notation and idea of [Mar18, Example 7.10].
Following [Rém02, Chapter 7] or [Mar18, Definition 7. 13], one can define a Kac–Moody algebra $\mathfrak{g}_{\mathcal{D}}$ of type $\mathcal{D}$ as a Lie algebra with generators $\mathfrak{h}_{\mathcal{D}}$, $\{e_{i}\}_{i\in I}$ and $\{f_{i}\}_{i\in I}$, where $e_{i}$ and $f_{i}$ are to be symbols satisfying the following relations.
$[\mathfrak{h}_{\mathcal{D}},\mathfrak{h}_{\mathcal{D}}]$ $=0$
$[h,e_{i}]$ $=c_{i}(h)e_{i}$
$[h,f_{i}]$ $=-c_{i}(h)f_{i}$
$[e_{i},f_{j}]$ $=-\delta^{i}_{j}h_{i}$
$(\operatorname{ad}e_{i})^{1-a_{ij}}e_{j}$ $=(\operatorname{ad}f_{i})^{1-a_{ij}}f_{j}=0\;(i\neq j)$
for $h\in\mathfrak{h}_{\mathcal{D}}$ and $i,j\in I$. In the following, we denote with $\mathfrak{g}(\mathbb{A})$ the Kac–Moody algebra of type $\mathcal{D}_{\mathrm{Kac}}^{\mathbb{A}}$.
Convention 3.4
For the rest of the article we consider the root datum $\mathcal{D}_{\mathrm{Kac}}^{\mathbb{A}}$, where $\mathbb{A}$ is a indecomposable, symmetrizable generalized Cartan matrix of size $n$ with rank $l$.
Let $\mathbb{A}$ be a generalized Cartan matrix and consider $\Pi=\{\alpha_{i}\mid i\in I\}$, then define the root lattice as
$Q\coloneqq\bigoplus_{i\in I}\mathbb{Z}\alpha_{i}$ (2)
and call an element $\alpha\neq 0\in Q$ a root if
$\{x\in\mathfrak{g}(\mathbb{A})\mid\operatorname{ad}_{h}(x)=\alpha(h)x\;\forall h\in\mathfrak{h}\}\neq\{0\}.$
Denote by $\Delta$ all roots of $\mathfrak{g}(\mathbb{A})$. Furthermore, a root $\alpha$ is called positive if the coefficients in $Q$ are non-negative, and negative otherwise. By [Mar18, Proposition 3.14 (5)] follows that every root is either positive or negative.
Next, recall the Weyl group of $\mathfrak{g}(\mathbb{A})$ from [Kac90, §3.7], [Mar18, Chapter 4.2] or [Rém02, 7.1.3]: it is the group $\mathcal{W}<\operatorname{GL}(\mathfrak{h})$ generated by the set $S=\{r_{1}^{\vee},\ldots,r_{n}^{\vee}\}$ of reflections
$r_{i}^{\vee}\colon\mathfrak{h}\to\mathfrak{h}\,,\;h\mapsto h-\alpha_{i}(h)\alpha_{i}^{\vee}.$
A root $\alpha\in\Delta$ is called real, if there exists a $w\in\mathcal{W}$, such that $w\alpha\in\Pi$. Denote by $\Delta^{\mathrm{re}}\coloneqq\mathcal{W}.\Pi$ the set of all real roots in $\mathfrak{g}(\mathbb{A})$.
3.2 The Tits Functor
Let $R$ be a ring and recall that $\mathcal{D}_{\mathrm{Kac}}^{\mathbb{A}}=\left(I,\mathbb{A},\Lambda_{\mathrm{Kac}},(c_{i})_{i\in I},(h_{i})_{i\in I}\right)$. Then a split torus scheme (cp. [Mar18, Section 7.4]) is given by
$T\colon\mathbb{Z}\text{-alg}\to\textbf{Grp}\,,\;T(R)\coloneqq\Lambda_{\mathrm{Kac}}^{\vee}\otimes_{\mathbb{Z}}R^{\times},$
alternatively, one can consider $T(R)$ as $\operatorname{Hom}_{\mathrm{grp}}(\Lambda_{\mathrm{Kac}},R^{\times})$:
$\Lambda^{\vee}\otimes R^{\times}\to\operatorname{Hom}_{\mathrm{grp}}(\Lambda_{\mathrm{Kac}},R^{\times})\,,\;h\otimes r\mapsto\left(\lambda\mapsto r^{\lambda(h)}\right).$
Since $\{v_{i}\}$ is a $\mathbb{Z}$-basis of $\Lambda_{\mathrm{Kac}}^{\vee}$ (see 3.3) and the rank is equal $2n-l$, the set
[…]
generates $T(R)$, cf. [Mar18, Example 7.55]. If $\mathbb{K}\in\{\mathbb{R},\mathbb{C}\}$, one can define an exponential map between $\mathfrak{h}$ and $T$ as follows
$\exp_{\operatorname{KM}}\colon\mathfrak{h}\to T(\mathbb{K})\,,\;(v_{i}\otimes r)\mapsto(v_{i}\otimes\exp(r)),$
where $\exp\colon\mathbb{K}\to\mathbb{K}^{\times}$ denotes the usual exponential map $\exp:(\mathbb{R},+)\to(\mathbb{R}^{\times},\cdot)$ or $\exp:(\mathbb{C},+)\to(\mathbb{C}^{\times},\cdot)$. Since $T(\mathbb{K})$ is isomorphic to $(\mathbb{K}^{\times})^{2n-l}$ and $\mathfrak{h}$ is isomorphic to $\mathbb{K}^{2n-l}$, the exponential map $\exp_{\operatorname{KM}}$ corresponds to the natural exponential map $\exp\colon\mathbb{K}^{2n-l}\to(\mathbb{K}^{\times})^{2n-l}$.
[…]
  1. KMG.1
    If $\mathbb{K}$ is a field, $\mathcal{G}(\mathbb{K})$ is generated by $\varphi_{i}(\operatorname{SL}_{2}(\mathbb{K}))$ and by $\eta(T(\mathbb{K}))$.
  2. KMG.2
    For every ring $R$, the homomorphism $\eta\colon\mathfrak{T}_{\Lambda}(R)\to\mathcal{G}(R)$ is injective.
  3. KMG.3
    For $i\in I$ and $r\in R^{\times}$, one has
    $\varphi_{i}\begin{pmatrix}r&0\\ 0&r^{-1}\end{pmatrix}=\eta\left(r^{h_{i}}\right),$
    where $r^{h_{i}}$ denotes the element $\lambda\mapsto r^{\langle\lambda,h_{i}\rangle}$ of the torus for each $h_{i}\in\Lambda^{\vee}$ and $\lambda\in\Lambda$.
  4. KMG.4
    If $\iota\colon R\to\mathbb{K}$ is an injective morphism of a ring $R$ in a field $\mathbb{K}$, then $\mathcal{G}(\iota)\colon\mathcal{G}(R)\to\mathcal{G}(\mathbb{K})$ is injective.
  5. KMG.5
    There is a homomorphism
    $\operatorname{Ad}\colon\mathcal{G}(\mathbb{C})\to\operatorname{Aut}([\mathfrak{g}(\mathbb{A}),\mathfrak{g}(\mathbb{A})])$
    whose kernel is contained in $\eta(\mathfrak{T}_{\Lambda}(\mathbb{C}))$, such that for $c\in\mathbb{C}$ and $i\in I$,
    $\operatorname{Ad}\left(\varphi_{i}\begin{pmatrix}1&c\\ 0&1\end{pmatrix}\right)=\exp(\operatorname{ad}ce_{i}),\;\operatorname{Ad}\left(\varphi_{i}\begin{pmatrix}1&0\\ c&1\end{pmatrix}\right)=\exp(\operatorname{ad}-cf_{i}),$
    and for $t\in\mathfrak{T}_{\Lambda}(\mathbb{C})$ and $i\in I$,
    $\operatorname{Ad}(\eta(t))(e_{i})=t(c_{i})\cdot e_{i},\;\operatorname{Ad}(\eta(t))(f_{i})=t(-c_{i})\cdot f_{i}.$
Note, that on fields, the Tits functor is unique up to isomorphism, see [Mar18, Theorem 7.82].
Convention 3.5
For the remainder of this article, let $\mathbb{K}\in{\mathbb{R},\mathbb{C}}$ and denote by $G$ the evaluation of the Tits functor $\mathcal{G}$ with respect to the Kac–Moody root datum $\mathcal{D}_{\text{Kac}}^{\mathbb{A}}$ on the field $\mathbb{K}$; $G$ is called a split (minimal) Kac–Moody group over $\mathbb{K}$ (of type $\mathbb{A}$) and $T$ its (standard) torus of $G$.
For any simple real root $\alpha_{i}\in\Delta^{re}$, $i\in I$, define the following subgroups of $G$ using the corresponding morphisms $\varphi_{i}$ from the definition of the Tits functor.
$U_{\alpha_{i}}\coloneqq\varphi_{i}\left(\left\{\begin{pmatrix}1&r\\ 0&1\end{pmatrix}\middle|\;r\in\mathbb{K}\right\}\right)\;\text{ and }\;\;U_{-\alpha_{i}}\coloneqq\varphi_{i}\left(\left\{\begin{pmatrix}1&0\\ -r&1\end{pmatrix}\middle|\;r\in\mathbb{K}\right\}\right).$
Now define the following group element for each simple root $\alpha_{i}$:
$\tilde{s}_{i}\coloneqq\varphi_{i}\begin{pmatrix}0&1\\ -1&0\end{pmatrix}.$
The group generated by these elements, $\widetilde{\mathcal{W}}\coloneqq\langle\tilde{s}_{i}\mid i\in I\rangle$, is called the extended Weyl group. A detailed discussion of the extended Weyl group and its properties can be found in [GHKW17, Section 18], [FHHK20, Section 3.13] or [KP85, p. 172]. Note that there is a canonical surjection from the extended Weyl group to the Weyl group given by
$\widetilde{\mathcal{W}}\to\mathcal{W}\,,\;\tilde{s}_{i}\mapsto r_{i}^{\vee}.$ (3)
The kernel of the map is given by $\widetilde{\mathcal{W}}\cap T$, cf. [KP85, Proposition 2.1] or [FHHK20, Section 3.13].
By the definition of a real root, each $\alpha\in\Delta^{re}$ can be written as $\alpha=w.\alpha_{i}$ for $\alpha_{i}\in\Pi$ and $w\in\mathcal{W}$. This translates, using the extended Weyl group, to the fact that for each real root $\alpha\in\Delta^{re}$ we can define the corresponding root (sub-)group $U_{\alpha}$ as follows
$U_{\alpha}=U_{w.\alpha_{i}}\coloneqq\tilde{w}U_{\alpha_{i}}\tilde{w}^{-1},$
where $\tilde{w}\in\widetilde{\mathcal{W}}$, see [FHHK20, Section 3.13]. Let $\alpha_{i}\in\Pi$, then one can describe an element of the root subgroup $U_{\alpha_{i}}$ as
$x_{\alpha_{i}}(r)\coloneqq\varphi_{i}\begin{pmatrix}1&r\\ 0&1\end{pmatrix}\in U_{\alpha_{i}},$
and an element of $U_{-\alpha_{i}}$ as
$x_{-\alpha_{i}}(r)\coloneqq\varphi_{i}\begin{pmatrix}1&0\\ -r&1\end{pmatrix}\in U_{-\alpha_{i}},$
for $r\in\mathbb{K}$.
Recall from Equation 2 that a root $\alpha$ is either positive or negative. Therefore, we can define $U_{+}=\langle U_{\alpha}\mid\alpha\text{ is positive}\rangle$ as the subgroup only generated by the positive root subgroups. On the same way we define $U_{-}$ as the subgroup of $G$ only generated by the negative root subgroups.
Further define for $\alpha\in\Delta^{\mathrm{re}}$ the corresponding rank one subgroup of $G$ as $G_{\alpha}\coloneqq\langle U_{\alpha},U_{-\alpha}\rangle$. Note that $G_{\alpha}\cong\operatorname{SL}_{2}(\mathbb{K})$, see [FHHK20, Proposition 3.17].
Define the (standard) Borel subgroups of $G$ by $B_{\pm}\coloneqq T\ltimes U_{\pm}$ (cp. [AB08, Corollary 7.122]) and denote all their conjugates as Borel subgroups.
Convention 3.6
In this article, we only consider Tits functors that are non-degenerate, i.e. we assume that the morphisms $\varphi_{i}$, $i\in I$, from the definition of the Tits functor do not map $\operatorname{SL}_{2}(\mathbb{K})$, where $\mathbb{K}$ is an arbitrary field, to the image of $U_{+}$ in the group given by the constructive Tits functor. The group given by the constructive Tits functor (cf. [Mar18, Definition 7.47]) is a Kac–Moody group defined by generators and relations.
The assumption of non-degeneracy allows us to apply [Mar18, Theorem 7.82 (2)]: Over fields, there is an isomorphism between the group given by the Tits functor and the group given by the constructive Tits functor. For details and a precise definition, see [Mar18, Definition 7.83] and [Mar18, Definition 7.47].
For a split Kac–Moody group $G$ over $\mathbb{K}$ of type $\mathbb{A}$ one defines the Kac–Peterson topology as the final group topology with respect to the maps $\varphi_{i}\colon\operatorname{SL}_{2}(\mathbb{K})\to G$ and $\eta\colon T\to G$, where $\operatorname{SL}_{2}(\mathbb{K})$ and $T$ are equipped with their natural Lie group topologies. For details of construction, see [HKM13, Section 7.5].
This topology turns $G$ into a topological group with the following properties:
  • $G$ is Hausdorff ([HKM13, Proposition 7.21] and [HK23, Definition and Remark 2.1]).
  • $G$ is a $k_{\omega}$-group ([HKM13, Proposition 7.10] and [HK23, Definition and Remark 2.1]).
  • The Kac–Peterson topology induces the unique Lie group topology on $T$, in particular $T$ is a closed subgroup ([HK23, Definition and Remark 2.1] and [HKM13, Corollary 7.17]).
For more details, refer to [HKM13, Chapter 7], [HK23] and [Har20]. Also, many of these properties were announced by Kac and Peterson in [KP83, Section 4G] without proof.
Now we prove theorem [FHHK20, Proposition 3.10] in the context of this article. This allows us to provide the torus with a Lie group topology.
Proposition 3.7
Let $G$ be a split Kac–Moody group over $\mathbb{K}$ of type $\mathbb{A}$, where $\mathbb{A}$ is an indecomposable, symmetrizable generalized Cartan matrix of size $n$ and rank $l$. Further denote by $\mathfrak{g}(\mathbb{A})$ the corresponding Kac–Moody algebra. Equip $\mathfrak{h}\cong\mathbb{K}^{2n-l}$ with the vector space topology and $T\cong(\mathbb{K}^{\times})^{2n-l}$ with the Kac–Peterson topology. Then the exponential map
$\exp_{\operatorname{KM}}\colon\mathfrak{h}\to T$
is a quotient map. In particular $T$ is a Lie group homeomorphic to $(\mathbb{K}^{\times})^{2n-l}$ equipped with its standard topology.
Proof:
The proof follows the same strategy as the proof of [FHHK20, Proposition 3.10], namely to consider $G$ as a subgroup of a larger Kac–Moody group with trivial center and to apply [HKM13, Proposition 7.18].
Recall that the root datum of $G$ is given by
$\mathcal{D}_{\mathrm{Kac}}^{\mathbb{A}}=(I,\mathbb{A},\Lambda_{\mathrm{Kac}},(c_{i})_{i\in I},(h_{i})_{i\in I}),$
where a $\mathbb{Z}$-basis for $\Lambda_{\mathrm{Kac}}$ is given by $\{u_{j}\}_{1\leq j\leq 2n-l}$ and for $\Lambda_{\mathrm{Kac}}^{\vee}$ by $\{v_{j}\}_{1\leq j\leq 2n-l}$, such that $h_{i}=v_{i}$ for $1\leq i\leq n$.
Now let $\mathbb{B}\in\mathbb{Z}^{2n-l\times 2n-l}$ be a generalized Cartan matrix with rank $2n-l$, such that $\mathbb{A}$ is included in $\mathbb{B}$ as a principal submatrix. This can be done in a similar way to expanding a non-invertible generalized Cartan matrix to obtain the extended Cartan algebra, see [Kac90, Proposition 1.1] or [Mar18, Section 3.5] (or 3.3). Consider now the corresponding root datum
$\widetilde{\mathcal{D}}=(J,\mathbb{B},\Lambda_{\mathrm{Kac}},(u_{i})_{i\in J},(v_{i})_{i\in J}),$
where $J=\{1,\ldots,2n-l\}$ and by construction follows that $\langle u_{j},v_{i}\rangle=b_{ij}$ and that $\langle u_{j},v_{i}\rangle=a_{ji}$ for $i,j\in I$. Denote by $\widetilde{G}$ the Kac–Moody group associated to $\widetilde{\mathcal{D}}$.
One can construct an embedding of $G$ into $\widetilde{G}$ as follows: For any $k\in\mathbb{K}$, $r\in\mathbb{K}^{\times}$ the embedding is given by
$\pi\colon G$ $\to\widetilde{G}$
$x_{\pm\alpha_{i}}(k)$ $\mapsto x_{\pm\alpha_{i}}(k)\;\;(1\leq i\leq n)$
$r^{v_{i}}$ $\mapsto r^{v_{i}}\;\;(i\in J),$
where $\alpha_{i}=c_{i}\otimes 1$ for $i\in I$. Compare this to [HK23, Proposition B.5] or [Voc23, Appendix C], where it is done for arbitrary split Kac–Moody groups.
Both are Kac–Moody groups and since we assumed that the Tits functor is non-degenerate, it is already clear that the relations of a Kac–Moody group (see [Mar18, Definition 7.47]) are preserved under the mapping $\pi$. Moreover, both groups can be equipped with the Kac–Peterson topology, and by construction is $G$ a subgroup of $\widetilde{G}$. Based on the construction, $\mathbb{B}$ is invertible and thus $\widetilde{\mathcal{D}}$ has finite center. This allows us to apply [HKM13, Proposition 7.18]. It follows from this proposition that the torus of $\widetilde{G}$ endowed with the Kac–Peterson topology is homeomorphic to $(\mathbb{K}^{\times})^{2n-l}$ equipped with its standard topology. Moreover the torus of $\widetilde{G}$ equals the torus of $G$ by construction of the root data. ∎
3.3 Twin Buildings
This section provides a short overview of the basic notions of a (twin) building. For more details see [AB08]. Let
$\mathcal{W}\coloneqq\langle r_{1},\ldots,r_{n}\mid(r_{i}r_{j})^{m_{ij}}=e\rangle,$
be a Coxeter group, where $(m_{ij})_{1\leq i,j\leq n}\subset\mathbb{Z}_{>0}\cup\{\infty\}$ such that
  • $m_{ii}=1$,
  • $m_{ij}=m_{ji}\geq 2$ for $i\neq j$.
The pair $(\mathcal{W},S)$ with generating set $S=\{r_{1},...,r_{n}\}$ is a Coxeter system. Note that any Weyl group of a Kac–Moody algebra forms a Coxeter system and the order of the elements are in relation with the entries of the generalized Cartan matrix, see [Mar18, Proposition 4.22].
Definition 3.8 (cp. Definition 5.1 from [AB08])
Given a Coxeter system $(\mathcal{W},S)$, a building $\Delta$ of type $(\mathcal{W},S)$ is a pair $(\mathcal{C},\delta)$, where $\mathcal{C}$ is a non-empty set whose elements are called chambers together with a Weyl distance function $\delta\colon\mathcal{C}\times\mathcal{C}\to\mathcal{W}$ such that for all $C,D,E\in\mathcal{C}$ the following properties hold:
  • $\delta(C,D)=e\Leftrightarrow C=D$;
  • if $\delta(C,D)=w\in\mathcal{W}$ and $\delta(E,C)=s\in S$, then $\delta(E,D)\in\{sw,w\}$. Further, if $l(sw)=l(w)+1$, then $\delta(E,D)=sw$, where $l(.)$ denotes the length of the reduced word in $\mathcal{W}$;
  • if $\delta(C,D)=w\in\mathcal{W}$, then for any $s\in S$ there exists a chamber $F\in\mathcal{C}$ such that $\delta(F,C)=s$ and $\delta(F,D)=sw$.
In order to associate a building to a group, we need the notion of a $BN$-pair, which implies the Bruhat decomposition. An important decomposition in order to define the structure of a building on a group, see [AB08, Section 6.2.6].
Now let $G$ be a group with subgroup $B$, $(\mathcal{W},S)$ a Coxeter system and $C\colon\mathcal{W}\to B\backslash G/B$ the bijection given by
$C(w)=B\tilde{w}B,$ (4)
where $\tilde{w}$ is a representative of the preimage of $w$ in the extended Weyl group under the surjection $\widetilde{\mathcal{W}}\to\mathcal{W}$, see Equation 3. This explicit description allows us to define a decomposition, as we will see later.
Next assume that for all $s\in S$ and $w\in\mathcal{W}$ the property $C(sw)\subseteq C(s)C(w)\subseteq C(sw)\cup C(w)$ is satisfied. This property allows us to define a building using the definition from above: consider the coset $G/B$ as the set of chambers and define the Weyl distance as follows
$\delta\colon G/B\times G/B$ $\to B\backslash G/B\to\mathcal{W}$
$(gB,hB)$ $\mapsto Bg^{-1}hB\mapsto C^{-1}(Bg^{-1}hB).$
This leads to the following equivalence, which allows to compute the Weyl distance
$\delta(gB,hB)=w\Longleftrightarrow g^{-1}h\in C(w).$
This construction satisfies the definition from above and thus leads to a building. In the following we denote by $\Delta_{G}$ a building which is associated to a group $G$.
The notion of a $BN$-pair guarantees the existence of a Bruhat decomposition and, thus, a building to act on.
Definition 3.9 (cp. Definition 6.55 from [AB08])
Let $G$ be a group with subgroups $B,N$. The pair $(B,N)$ is called $BN$-pair for $G$ if the following is satisfied:
  • $G$ is generated by $B$ and $N$.
  • The intersection $N\cap B$ is normal in $N$.
  • The quotient $N/T$ admits a set of generating involutions $S$ such that
    • for all $w\in N/T$ and $s\in S$ is $sBw\subseteq BswB\cup BwB$, and
    • $sBs^{-1}$ is no subgroup of $B$ for all $s\in S$.
We denote by $T$ the intersection $N\cap B$ and by $\mathcal{W}$ the quotient $N/T$. That a group $G$ has a $BN$-pair $(B,N)$ implies the Bruhat decomposition
$G=\bigsqcup_{w\in\mathcal{W}}BwB,$
see [AB08, Section 6.2.6] By this decomposition and using Equation 4 follows that there is the desired bijection
$C\colon\mathcal{W}$ $\to B\backslash G/B$
$w$ $\mapsto C(w)\coloneqq BwB.$
Given two buildings $\Delta_{+}=(\mathcal{C}_{+},\delta_{+})$ and $\Delta_{-}=(\mathcal{C}_{-},\delta_{-})$ of type $(\mathcal{W},S)$, where $\mathcal{C}_{-}\cap\mathcal{C}_{+}=\emptyset$. Then call the triple
$\Delta_{\pm}\coloneqq(\mathcal{C}_{+},\mathcal{C}_{-},\delta^{\star}),$
where the map
$b\delta^{\star}\colon(\mathcal{C}_{+}\times\mathcal{C}_{-})\cup(\mathcal{C}_{-}\times\mathcal{C}_{+})\to\mathcal{W}$
satisfies for any $C\in\mathcal{C}_{\varepsilon}$, $D\in\mathcal{C}_{-\varepsilon}$ and $\varepsilon\in\{+,-\}$
  • $\delta^{\star}(C,D)=\delta^{\star}(D,C)^{-1}$.
  • If $\delta^{\star}(C,D)=w\in\mathcal{W}$ and if $\delta_{\varepsilon}(E,C)=s\in S$ for $E\in\mathcal{C}_{\varepsilon}$ and $\langle(sw)<l(w)$, then $\delta^{\star}(E,D)=sw$.
  • For any $s\in S$ there exists a chamber $E\in\mathcal{C}_{\varepsilon}$ such that $\delta_{\varepsilon}(E,C)=s$ and $\delta^{\star}(E,D)=sw$.
a twin building of type $(\mathcal{W},S)$ with codistance function $\delta^{\star}$. For more details, see [AB08, Definition 5.133].
By [AB08, Definition 6.78] can a building constructed from a group $G$ if it has a twin $BN$-pair $(B_{+},B_{-},N)$, which is a $BN$-pair with respect to the subgroup $B_{+}$ and $B_{-}$. Similar to the case for one building, a twin $BN$-pair guaranties the Birkhoff decomposition (cf. [AB08, Proposition 6.81]):
$G=\bigsqcup_{w\in\mathcal{W}}B_{\varepsilon}wB_{-\varepsilon},$
where $\varepsilon\in\{+,-\}$. Again, as in the case for one building, this decomposition allows to define a codistance function on the set of chambers $G/B_{\pm}$. Note that $B_{+},B_{-}$ and $N$ are subgroups of $G$ satisfying some properties, for details see [AB08, Definition 6.78].
Examples of groups that possess a (twin) $BN$-pair are the general linear group ([AB08, Chapter 6.5]), the symplectic group ([AB08, Chapter 6.6]), orthogonal groups ([AB08, Chapter 6.7]) and unitary groups ([AB08, Chapter 6.8]).
Remark 3.10
Since we are interested in Kac–Moody groups, the notion of an RGD-system is useful, see [AB08, Section 8.6.1] and [AB08, Definition 7.82] for details. An RGD system allows to construct a $BN$-pair and thus associate a building to a group.
Let $G$ be a Kac–Moody group over $\mathbb{K}$ of type $\mathbb{A}$. Then, by [Mar18, Theorem 7.69], the triple $(G,(U_{\alpha})_{\alpha\in\Delta^{\mathrm{re}}},T)$ represents a (linear) RGD system. This means that this triple satisfies the following properties
  1. RGD.1
    For all $\alpha\in\Delta^{\mathrm{re}}$ the root group $U_{\alpha}$ is non-trivial.
  2. RGD.2
    For all $\alpha,\beta\in\Delta^{\mathrm{re}}$ with $\alpha\neq\beta$ such that the set $\{\alpha,\beta\}$ is prenilpotent (see [AB08, Definition 8.41]), follows that $[U_{\alpha},U_{\beta}]\leq U_{(\alpha,\beta)}$, where $(\alpha,\beta)\coloneqq[\alpha,\beta]\setminus\{\alpha,\beta\}$.
  3. RGD.3
    For every $s\in S$ there exists a function $m\colon U_{\alpha_{s}}\setminus\{e\}\to G$, such that for all $u\in U_{\alpha_{s}}\setminus\{e\}$ and $\alpha\in\Delta^{\mathrm{re}}$ follows that
    $m(u)\in U_{-\alpha_{s}}uU_{-\alpha_{s}}$
    $m(u)U_{\alpha}m(u)^{-1}=U_{s\alpha}.$
    Furthermore, for all $u,v\in U_{\alpha_{s}}\setminus\{e\}$ follows $m(u)^{-1}m(v)\in T$.
  4. RGD.4
    For all $s\in S$ is $U_{-\alpha_{s}}$ no subgroup of $U_{+}$.
  5. RGD.5
    $G=T\langle U_{\alpha}\mid\alpha\in\Delta^{\mathrm{re}}\rangle$
  6. RGD.6
    $T\leq\bigcap_{\alpha\in\Delta^{\mathrm{re}}}N_{G}(U_{\alpha})$
Following [AB08, Chapter 8.8] one can define subgroups $B_{\pm}$ by
$B_{\pm}\coloneqq TU_{\pm},$
see [AB08, Remark 7.87]. Further in [AB08, Chapter 8.8] it is shown, that $N_{G}(U_{\pm})=B_{\pm}$ and that $B_{+}\cap B_{-}=T$ ([AB08, Corollary 8.78]).
All together, it is shown in [AB08, Theorem 8.80] that an RGD system leads to an $BN$-pair with bijection
$N_{G}(T)/T\cong\mathcal{W}.$
By [AB08, Corollary 8.79] follows that the last RGD axiom is an equality,
$T=\bigcap_{\alpha\in\Delta^{\mathrm{re}}}N_{G}(U_{\alpha}).$
3.4 A Characteristic Subgroup
A very useful subgroup to study symmetric spaces which belongs to a Kac–Moody group is the centered subgroup, that is the subgroup only generated by the root groups.
Definition 3.11
Let $G$ be a split Kac–Moody group over $\mathbb{K}$ of type $\mathbb{A}$. Denote with $G^{\dagger}<G$ the subgroup of $G$ generated by the root subgroups of $G$, i.e.
$G^{\dagger}\coloneqq\langle U_{\alpha}\mid\alpha\in\Delta^{re}\rangle.$
The group is called a centered subgroup.
Note, that from 3.10 it is clear, that $G^{\dagger}$ is a normal subgroup of $G$. From [AB08, Corollary 8.79] it follows that the torus of a Kac–Moody group can be described as the intersection of all normalizers of the root groups. We use this fact to define a torus for $G^{\dagger}$:
$T^{\dagger}\coloneqq\bigcap_{\alpha\in\Delta^{\mathrm{re}}}N_{G^{\dagger}}(U_{\alpha})=T\cap G^{\dagger}.$
Since the RGD axioms are manly using root groups, it follows that $(G^{\dagger},(U_{\alpha})_{\alpha\in\Delta^{\mathrm{re}}},T^{\dagger})$ forms an RGD system. Moreover, according to [AB08, Proposition 8.82], the kernel of the action of $G$ on the associated twin building $\Delta_{\pm}$ is contained in the torus $T$, so that
$N_{G^{\dagger}}(T^{\dagger})/T^{\dagger}\cong\mathcal{W}.$
Hence the type of the twin building associated to $G^{\dagger}$ is the same as the type of the twin building associated to $G$.
Lemma 3.12
Let $G$ be a split Kac–Moody group over $\mathbb{K}$ of type $\mathbb{A}$. Then $G^{\dagger}$ is the derived subgroup of $G$, in particular a characteristic subgroup, i.e.
$f(G^{\dagger})\subseteq G^{\dagger}$
for all $f\in\operatorname{Aut}(G)$.
Proof:
By the RGD axioms one has $G=TG^{\dagger}$ (see [AB08, Chapter 8]). Moreover, $G^{\dagger}$ is a normal subgroup of $G$, since
$T=\bigcap_{\alpha\in\Delta^{re}}{N_{G}(U_{\alpha})},$
cp. [AB08, Corollary 8.79]. Furthermore $G/G^{\dagger}$ is abelian, therefore $[G,G]\subseteq G^{\dagger}$.
On the other hand, by definition $G^{\dagger}$ is generated by the root subgroups, in particular $G^{\dagger}$ is generated by the standard rank one subgroups
$G_{i}=\langle U_{\alpha_{i}},U_{-\alpha_{i}}\rangle\cong\operatorname{SL}(2,k).$
Due to the fact that $\operatorname{SL}(2,k)$ is perfect for $k\in\{\mathbb{R},\mathbb{C}\}$, one deduces that $G^{\dagger}$ is perfect. Thus
$[G,G]\subseteq G^{\dagger}=[G^{\dagger},G^{\dagger}]\subseteq[G,G]$
which finishes the proof. ∎
Note, that there is also an action of $\operatorname{Aut}(G)$ on $G^{\dagger}$ given by the homomorphism $\operatorname{Aut}(G)\to\operatorname{Aut}(G^{\dagger})$ which restricts any $f\in\operatorname{Aut}(G)$ on $G^{\dagger}$.
Remark 3.13
Since we are interested in a generalization of [FHHK20], it is useful to know that the centered subgroup $G^{\dagger}$ corresponds to the type of Kac-Moody group used by the authors in [FHHK20] to develop the theory of symmetric Kac–Moody spaces. In particular, the type of Kac–Moody groups used in [FHHK20] is called centered, which means that the group is generated only by its root subgroups, without an additional torus. In the language of root data, the group is generated by a simply connected root datum, cf. [Mar18, Example 7.11].
3.5 The Chevalley Involution
The first goal of this article is to construct a symmetric space with respect to a Kac–Moody group. For this purpose, it is useful and important to define an involution. Therefore let $G$ be a Kac–Moody group over $\mathbb{K}\in\{\mathbb{R},\mathbb{C}\}$ of type $\mathbb{A}$. Then $\Theta\colon G\to G$ denotes a continuous involution which is given for $\mathbb{K}=\mathbb{C}$ by
$\Theta(x_{\alpha}(r))=x_{-\alpha}(\overline{r})\,,\;\Theta(r^{v_{i}})=\overline{r}^{-v_{i}},$
for $r\in\mathbb{C}$ and $\alpha\in\Delta^{\mathrm{re}}$. Since the complex conjugation acts trivially on the reals, it can be omitted in the case $\mathbb{K}=\mathbb{R}$. In [Cap09, §8.2] it is described how this involution is obtained by lifting the Chevalley involution of a Kac–Moody algebra $g(\mathbb{A})$ (cf. [Kac90, §1.3]) to a continuous involution on $G$. In particular, this means that $\Theta$ maps $U_{\alpha}$ to $U_{-\alpha}$ and stabilizes the torus $T$.
Denote by $K=G^{\Theta}$ the set of fixed points of $\Theta$ and define by $\tau\colon G\to G$, $g\mapsto g\Theta(g)^{-1}$ the twist map of $G$. In the following lemma, we define useful subgroups and analyze their properties. The statements are based on [FHHK20, Lemma 3.24] and [FHHK20, Lemma 3.26].
Lemma 3.14
Let $G$ be a split Kac–Moody group over $\mathbb{K}$ of type $\mathbb{A}$, where $\mathbb{K}\in\{\mathbb{R},\mathbb{C}\}$. Define the subgroups $A\coloneqq\tau(T)$ and $M\coloneqq K\cap T$. Then it follows that $T\cong M\times A$ and the following properties are satisfied
  1. (A1)
    For $g,h\in G$ one has $hK=gK\Leftrightarrow\tau(g)=\tau(h)$ and $\tau^{-1}(e)=K$.
  2. (A2)
    The subgroup $A$ is isomorphic to $(\mathbb{R}_{>0})^{2n-l}$.
  3. (A3)
[…]
  4. (A4)
    $M\cap A=\{e\}=K\cap A$
  5. (A5)
    $K\cap B_{\pm}=M$.
Proof.
For (A1) let $g,h\in G$ and it follows from $gK=hK$ that there is a $k\in K$ such that $g=hK$. For one direction one calculates
$gK=hK\Rightarrow\tau(g)=\tau(hk)=\tau(h),$
and for the other direction
$\tau(g)=\tau(h)$ $\Rightarrow g\Theta(g)^{-1}=h\Theta(h)^{-1}$
$\Rightarrow\Theta(h^{-1}g)=h^{-1}g$
$\Rightarrow h^{-1}g\in K$
$\Rightarrow gK=hK.$
To see that $\tau^{-1}(e)=K$, let $g\in G$ be an element with $\tau(g)=e$. Then calculate
$\tau(g)=e\Leftrightarrow\Theta(g)=g\Leftrightarrow g\in K.$
In order to prove (A2), we first check, that for an element $t\in T$ holds
$\tau(t)=\begin{cases}t^{2}&\text{for }\mathbb{K}=\mathbb{R}\\ |t|^{2}&\text{for }\mathbb{K}=\mathbb{C}.\end{cases}$
This follows directly from the definition of the action of $\Theta$ on $T$, to be precise, for $t\in T$ applies $\Theta(t)=\overline{t}^{-1}$ for $\mathbb{K}=\mathbb{C}$ and $\Theta(t)=t^{-1}$ for $\mathbb{K}=\mathbb{R}$. Hence, it follows that $A=\tau(T)\cong(\mathbb{R}_{>0})^{2n-l}$, since the dimension of $T$ is $2n-l$.
Next, we want to check (A4). Let $g\in K\cap A\subseteq T$. Then by definition of $K$ and $\Theta$ it follows that
$g=\Theta(g)=\overline{g}^{-1}.$
Thus $g\overline{g}=e$ and from the fact that $A=\tau(T)\cong(\mathbb{R}_{>0})^{2n-l}$ it follows that $g=e$. In particular, the statement that $M\cap A=\{e\}$ is true follows then directly from $M\subset K$.
Finally, we prove the statement of (A5). Therefore recall, that by the definition of $B_{\pm}$ it follows that $\Theta(B_{\pm})=B_{\mp}$ and that $T=B_{+}\cap B_{-}$. Note, that this is based on the fact that $U_{+}\cap U_{-}=\{e\}$. Thus, it follows that $K\cap B_{\pm}=K\cap T$. Now let $g\in K\cap B_{\pm}$, then follows that $g\in T$. Using the decomposition $T\cong M\times A$ we can write $g=ma$ for $m\in M$ and $a\in A$., one calculates
$\tau(g)=\begin{cases}g^{2}=m^{2}a^{2}=a^{2}&\text{for }\mathbb{K}=\mathbb{R}\\ |g|^{2}=|m|^{2}|a|^{2}=|a|^{2}=a^{2}&\text{for }\mathbb{K}=\mathbb{C}.\end{cases}$
On the other hand, by the fact that $g\in K$ follows that $\tau(g)=e$. This leads to $a^{2}=e$ and since $a\in A\cong(\mathbb{R}_{>0})^{2n-l}$ it follows that $\tau(g)=e$ if and only if $a=e$.
The fact that $T\cong M\times A$ follows from the polar decomposition:
[…]
3.6 The Iwasawa Decomposition
In the theory of symmetric spaces, it is necessary to define flats because they are the most important principles for understanding actions and causal boundaries. To do so, one needs the topological Iwasawa decomposition, i.e. that the multiplication induces a homeomorphism
$m_{\pm}\colon U_{\pm}\times A\times K\to G.$
In general, this is known for centered Kac–Moody groups, i.e. Kac–Moody groups generated only by their root groups without any additional torus, see [KP85, Section 3].
First we will prove the algebraic statement that for a Kac–Moody group $G$ there is a decomposition $G=U_{\pm}AK$ which induces a bijection. After establishing the algebraic Iwasawa decomposition, we will use the same strategy as in the proof of [FHHK20, Theorem 3.31 (ii)] to prove the topological assertion for Kac–Moody groups over $\mathbb{K}$ of arbitrary type $\mathbb{A}$.
The following section is mainly based on [FHHK20, Section 3.30]. Denote by $\mathbb{K}$ the field of real or complex numbers and with $\mathbb{A}$ a symmetrizable, indecomposable generalized Cartan matrix.
Proposition 3.15
Let $G$ be a Kac–Moody group over $\mathbb{K}$ and of arbitrary type $\mathbb{A}$. Then $G$ decomposes into
$G=U_{\pm}AK=KAU_{\pm}.$
Moreover, the maps,
$m_{\pm}\colon U_{\pm}\times A\times K\to G$
given by multiplication, are bijective.
Proof:
By 3.10 and Section 3.4, one can decompose $G$ into $TG^{\dagger}$, where $T$ is the torus and $G^{\dagger}$ is the centered subgroup of $G$. Since $\mathbb{K}$ plays no role in the algebraic decomposition, and $G^{\dagger}$ is a centered Kac–Moody group, one can use [FHHK20, Theorem 3.31 (ii)]. Therefore denote by $K^{\dagger}$ the fixed point set $(G^{\dagger})^{\Theta}$ and by $B_{\pm}^{\dagger}$ the (standard) Borel subgroup $B_{\pm}^{\dagger}=T^{\dagger}\ltimes U_{\pm}$. Then there is the decomposition
$G^{\dagger}=B_{\pm}^{\dagger}K_{\pm}^{\dagger}.$
Recall that $T=MA$, $M\subseteq K$ by 3.14 and that $T$ normalizes $U_{\pm}$. Also denote with $A^{\dagger}=\tau(T^{\dagger})$ and with $M^{\dagger}=K\cap T$. All properties of 3.14 are also true for the corresponding subgroups in the context of $G^{\dagger}$. It follows
$G$ $=TG^{\dagger}$
$=TB_{\pm}^{\dagger}K^{\dagger}$
$=TU_{\pm}A^{\dagger}M^{\dagger}K^{\dagger}$
$=U_{\pm}(MA)A^{\dagger}K^{\dagger}=U_{\pm}AK$
since $K\setminus K^{\dagger}$ can by only contained in $T$ and by 3.14 (A5) it is contained in $M$. The other order of the decomposition results from the symmetry arguments.
To see that these maps $m_{\pm}$ are bijective, we need to check injectivity. To do this, we assume that $u_{1}a_{1}k_{1}=u_{2}a_{2}k_{2}$. By reordering follows,
$K\ni k_{1}k_{2}^{-1}=a_{1}^{-1}u_{1}^{-1}u_{2}a_{2}\in AU_{\pm}A\subseteq B_{\pm}.$
Hence $a_{1}^{-1}u_{1}^{-1}u_{2}a_{2}\in K$ and $B_{\pm}\cap K=M\subseteq T$, see 3.14. It follows that
$u_{1}^{-1}u_{2}\in U_{\pm}\cap T=\{1\}\Longrightarrow u_{1}=u_{2},$
(cp. [AB08, Lemma 7.62]) and from this, that
$a_{1}^{-1}a_{2}\in M\cap A=\{1\}\Longrightarrow a_{1}=a_{2}.$
It follows that $k_{1}=k_{2}$ and thus $m_{\pm}$ is injective. ∎
In the following we consider only Kac–Moody groups of non-spherical type $\mathbb{A}$. In the case where $\mathbb{A}$ is of spherical type, the Kac–Moody group of type $\mathbb{A}$ is a finite-dimensional Lie group and the topological Iwasawa decomposition is a well-known fact, see for example [Kna02, Theorem 6.46]. Further, there are also proofs of the topological Iwasawa decomposition with respect to other groups, e.g. for $\operatorname{GL}_{n}(\mathbb{R})$ ([PR94, Proposition 3.12]) or in greater generality for reductive $\mathbb{R}$-groups ([PR94, Theorem 3.9]).
Now for the topological version, the idea is to use the same strategy as in the proof of [FHHK20, Theorem 3.31 (ii)], to be precise we will construct the inverse functions and show that they are open. That we can use the same strategy as in [FHHK20] is based on the statement that every Kac–Moody group has a strong topological building, as shown by Grüning and the first-named author of this article in [HK23, Appendix B] for centered Kac–Moody groups. The very technical generalization for arbitrary split Kac–Moody groups can be found in [Voc23, Appendix C]. For an introduction to topological buildings see [HKM13, Section 3.1] and for strong topological buildings in particular see [HKM13, Definition 3.21].
Lemma 3.16
Let $G$ be a Kac–Moody group over $\mathbb{K}$ and of type $\mathbb{A}$, where $\mathbb{A}$ is non-spherical. Then the fibration
$p_{\pm}\colon\nicefrac{{G}}{{AU_{\pm}}}\to\nicefrac{{G}}{{B_{\pm}}}$
is trivial.
Proof:
First, note that $G$ leads to a non-spherical strong topological twin building, according to a result by Grüning and the first-named author of this article, see [HK23, Corollary B.8] (or [Voc23, Appendix C]). Furthermore, note that $T\cong A\times M$ as topological groups and that $U_{\mp}\times T\times U_{\pm}$ is open. In particular the map $U_{\mp}\times T\times U_{\pm}\to B_{\mp}B_{\pm}$ is a homeomorphism (cf. [AB08, Section 8.8] and [HK23, Corollary B.7]) and $B_{\mp}B_{\pm}$ is open (cf. [HKM13, Lemma 6.1]). Hence
$\{e\}\times M\times\{e\}\to U_{\mp}\times T\times U_{\pm}\to U_{\mp}\times A\times U_{\pm}$
provides a local trivialization of the fibration
$M\to\nicefrac{{G}}{{AU_{\pm}}}\to\nicefrac{{G}}{{B_{\pm}}},$
in particular, every point in $G/AU_{\pm}$ has a neighborhood that looks like an open set in $G/B_{\pm}$ times $M$.
Since the non-spherical twin building associated to $G$ is a strong topological twin building, it is contractible (cf. [HKM13, Theorem 5.13]) and thus the fibration $p_{\pm}$ is trivial. ∎
For the next lemma we need the notion of $k_{\omega}$-spaces: a Hausdorff topological space $X$ is called a $k_{\omega}$-space, if there is a family of compact subspaces $(X_{n})_{n\in\mathbb{N}}$, called $k_{\omega}$-sequence, such that
$X=\bigcup_{n\in\mathbb{N}}X_{n}$
and any open subset $Y\subseteq X$ is open in $X$ if and only if for every $n\in\mathbb{N}$ $Y\cap X_{n}$ is open in $X_{n}$. Call the pair $(X,(X_{n})_{n\in\mathbb{N}})$ $k_{\omega}$-pair. For more details about $k_{\omega}$-spaces see [FT77] or [FHHK20, Definition 3.33 et seq.]. Here we recall two lemmas about $k_{\omega}$-spaces for the reader’s convenience that we use in order to derive the topological Iwasawa decomposition.
Lemma 3.17 (Lemma 3.34 from [FHHK20])
Let $(X,(X_{n}))$ and $(Y,(Y_{n}))$ be $k_{\omega}$ pairs and let $f\colon X\to Y$ be a continuous bijection such that for all $n\in\mathbb{N}$ there exists a $m\in\mathbb{N}$ such that $Y_{n}\subset f(X_{m})$.
Then $f(X_{n})$ is a $k_{\omega}$-sequence for $Y$ and $f$ is a homeomorphism.
Lemma 3.18 (Lemma 3.35 from [FHHK20])
Let $(X,(X_{n}))$ be a $k_{\omega}$ pair and let $\pi\colon\widetilde{X}\to X$ be a finite-sheeted covering. Then $(\widetilde{X},\pi^{-1}(X_{n}))$ is a $k_{\omega}$ pair.
Now, we can prove the following, which is a version of [FHHK20, Proposition 3.36] adapted to the context of this article.
Lemma 3.19
Let $G$ be a Kac–Moody group over $\mathbb{K}$ of type $\mathbb{A}$, where $\mathbb{A}$ is non-spherical and define the maps
$\iota_{\pm}\colon K$ $\to G/AU_{\pm}$
$k$ $\mapsto kAU_{\pm}.$
These maps are homeomorphisms.
Proof:
The goal is to use 3.17. According to the algebraic Iwasawa decomposition (cf. 3.15), the maps $\iota_{\pm}$ are bijections, since multiplication is continuous, these maps are also continuous.
Next, we recall that $G$ acts strongly transitively on the associated twin building (cf. 3.10) and we can therefore use the Birkhoff decomposition according to [AB08, Proposition 6.75]. Based on the Birkhoff decomposition we can define the following:
$G_{l}^{\pm}\coloneqq\bigcup_{w\in\mathcal{W},\,l(w)\leq l}B_{\pm}wB_{\pm},$
and denote with $\widetilde{D}_{l,\pm}$ the image of $G_{l}^{\pm}$ in $G/AU_{\pm}$ and with $D_{l,\pm}$ the image of $G_{l}^{\pm}$ in $G/B_{\pm}$. By [HKM13, Corollary 3.13] follows that
$\left(\nicefrac{{G}}{{B_{\pm}}},\left(D_{l,\pm}\right)\right)$
is a $k_{\omega}$-pair. Note that the preimage of $D_{l,\pm}$ under the fibration $p_{\pm}$ is exactly $\widetilde{D}_{l,\pm}$ and that the fibration by 3.16 is trivial, so we can apply 3.18, which works here, since the statement is just a topological assertion about $k_{\omega}$ pairs, so we observe that
$\left(\nicefrac{{G}}{{AU_{\pm}}},\left(\widetilde{D}_{l,\pm}\right)\right)$
is also a $k_{\omega}$-pair.
In order to apply 3.17, it remains to prove that $(K,(K_{l}^{\pm}))$ is a $k_{\omega}$-pair, where $K_{l}^{\pm}\coloneqq K\cap G_{l}^{\pm}$. Therefore by [HKM13, Corollary 7.11] one knows that
[…]
and in particular that
[…]
To conclude that $(K,(K_{l}^{\pm}))$ is a $k_{\omega}$ pair, we must show that the subsets $(K_{l}^{\pm})$ are compact. Therefore, recall that $G$ can be written as $KB_{\pm}$ using 3.15, and hence $K$ acts chamber-transitive on the to $G$ associated twin building $\Delta_{\pm}$. By the construction of $G_{l}^{\pm}$ and $K_{l}^{\pm}$ it follows that
$G_{l}^{\pm}=K_{l}^{\pm}(G_{l}^{\pm}\cap B_{\pm}).$
Denote by $\mathcal{C}_{\pm}$ the set of chambers of $\Delta_{\pm}$, then by using the Bruhat order ([AB08, Exercise 3.59 (c)]) we can define the set
$E_{w}(C)=\{D\in\mathcal{C}_{\pm}\mid\delta_{\pm}(C,D)=w\}\subseteq\Delta_{\pm},$
where $C\in\mathcal{C}_{\pm}$ denotes a chamber in $\Delta_{\pm}$, $w\in\mathcal{W}$ an element of the Weyl group and $\delta_{\pm}$ the Weyl distance in relation to the two halves of the twin building. By the previous discussion $K_{l}^{\pm}$ acts chamber-transitive on
$\bigcup_{w\in\mathcal{W}\,,\;l(w)\leq l}E_{w}(C).$
Moreover, this is a finite union of compact sets, see [HKM13, Corollary 3.10].
Using again the chamber-transitivity it follows that we can rewrite the union as the quotient
$K_{l}^{\pm}/(K_{l}^{\pm}\cap B_{\pm}).$
Recall that $M$ is the compact subgroup of $T$, in particular $M$ is a Lie group, and that $K_{l}^{\pm}\cap B_{\pm}\subseteq M$. Thus, by the discussion around [Pal61, Corollary 4.1], it follows that
$K\to K/(K\cap B_{\pm})$
are fibers. $K_{l}^{\pm}$ are closed sets so it follows that
$\pi\colon K_{l}^{\pm}\to K_{l}^{\pm}/(K_{l}^{\pm}\cap B_{\pm})$
are also fibers and the corresponding fiber bundles are compact. Thus $K_{l}^{\pm}$ is a compact set for every $l\in\mathbb{N}$, therefore, the pair $(K,(K_{l}^{\pm}))$ is a $k_{\omega}$ pair. Together with the fact that
$\iota_{\pm}(K_{l}^{\pm})=\widetilde{D}_{l,\pm}$
we can apply 3.17. ∎
At this point, we are able to prove the topological Iwasawa decomposition for Kac–Moody groups of arbitrary type over $\mathbb{C}$.
Theorem 3.20
Let $G$ be a split Kac–Moody group over $\mathbb{K}$ of type $\mathbb{A}$, where $\mathbb{A}$ is non-spherical. Then each of the multiplication maps
$m_{\pm}$ $\colon U_{\pm}\times A\times K\to G$
is a homeomorphism.
Proof:
By the algebraic Iwasawa decomposition 3.15 we know that $m_{\pm}$ is a bijection, so it follows from the continuity of multiplication in $G$ that $m_{\pm}$ is continuous with respect to the Kac–Peterson topology. Thus, all that remains is to construct an inverse continuous map as in the proof of [FHHK20, Proposition 3.31].
Let $g\in G$ and define the two elements
$k(g)\coloneqq\iota_{\pm}^{-1}\left(gAU_{\pm}\right)\;\text{ and }\;h(g)\coloneqq k(g)^{-1}g,$
where the maps $\iota_{\pm}$ are the same as in 3.19. Therefore we define the map
$n_{\pm}\colon G$ $\to K\times AU_{\pm}$
$g=k(g)h(g)$ $\mapsto(k(g),h(g)),$
which are by construction inverse to $m_{p}m$ and they are continuous, since $G$ is a topological group and $\iota_{\pm}$ is a homeomorphisms. It is left to check, that the multiplication maps $A\times U_{\pm}\to AU_{\pm}$ are open, but this follows from [HKM13, Proposition 7.29]. ∎
A useful application of the Iwasawa decomposition is the following lemma, which can be found for centered Kac–Moody groups over $\mathbb{R}$ in [FHHK20, Lemma 3.27].
Lemma 3.21
Let $G$ be a Kac–Moody group over $\mathbb{K}$ of type $\mathbb{A}$. Then $N_{G}(T)\cap\tau(G)=A$.
Proof.
The following applies: $\tau(t)=A$ and $T\subseteq N_{G}(T)$, i.e. $A\subseteq N_{G}(T)\cap\tau(G)$. For the converse implication let $g\in N_{G}(T)\cap\tau(G)$. By 3.15 there exists elements $u\in U_{+}$, $t\in T$ and $k\in K$, such that
$g=\tau(utk)=u\tau(t)\Theta(u)^{-1}\in U_{+}\tau(t)U_{-}.$
Since the RGD system obtained from a Kac–Moody group is refined (see [Rém02, Theorem 1.5.4]), we can use the refined Birkhoff decomposition
$G=\bigsqcup_{n\in N_{G}(T)}U_{+}nU_{-},$
see [Kum02, Theorem 5.2.3 (g)].
Since $g\in N_{G}(T)$ it follows that $g=\tau(t)$ and hence $g\in A$. ∎
4 Kac–Moody Symmetric Spaces
4.1 A criterion for symmetric spaces
In the following section we prove a statement in the context of buildings that allows us to show that $K\cap\tau(G)=\{e\}$. This result will be a very powerful statement to show that a Kac–Moody group can be transformed into a symmetric space in the sense of Loos. The following lemma is very technical and uses building theory arguments, for a brief introduction see Section 3.3. Let $\mathbb{K}\in\{\mathbb{R},\mathbb{C}\}$ and $\mathbb{A}$ be a symmetrizable, indecomposable generalized Cartan matrix.
In order to state the lemma, call an element $g\in G$ of a Kac–Moody group $G$ with the property
$\Theta(g)=g^{-1}$
symmetric.
Lemma 4.1 (Lemma 4.2 in [Hor17])
Let $G$ be a Kac–Moody group over $\mathbb{K}$ of type $\mathbb{A}$ and denote with $\Delta_{\pm}$ the associated twin building. For a symmetric element $g\in G$ the statements below are equivalent:
  1. 1
    The element $g$ fixes a $\Theta$-stable twin apartment chamberwise. $\Theta$-stable means that a twin apartment of $\Delta_{\pm}$ is invariant as a set under the action of $\Theta$.
  2. 2
    The element $g$ fixes a twin apartment chamberwise.
  3. 3
    The element $g$ stabilizes a chamber.
  4. 4
    For all chambers $D$ in the twin building $\Delta_{\pm}$, the length of any minimal gallery from $D$ to a chamber in the set of $G$-orbits $\{g^{n}.D\mid n\in\mathbb{Z}\}$, is bounded.
  5. 5
    For some chamber $D$ in the twin building $\Delta_{\pm}$, the length of any minimal gallery from $D$ to a chamber in the set of $G$-orbits $\{g^{n}.D\mid n\in\mathbb{Z}\}$, is bounded.
  6. 6
    The element $g$ stabilizes a spherical residue in either half of the twin building.
Remark 4.2
This statement was formulated and proven by Max Horn in [Hor17]. We reproduce the proof here and, in fact, add some details for the reader’s convenience.
Proof:
[…]
which gives us the desired statement.
[…]
    Let $D$ be a chamber in $\Delta_{\pm}$ which is stabilized by $g$, i.e. $g.D=D$. If $g$ stabilizes $D$, then also $g^{-1}$ stabilizes $D$:
    $D=(g^{-1}g).D=g^{-1}.(g.D)=g^{-1}.D.$
    Hence one obtains by a straight forward calculation,
    $\Theta(D)=\Theta(g.D)=g^{-1}.\Theta(D),$
    that $g$ stabilizes the $\Theta$-stable twin apartment $\Sigma(D,\Theta(D))$, since $\Theta(D)$ is opposite of $D$.
[…]
    First, recall that the building $\Delta_{\pm}$ associated with the Kac–Moody group $G$ is of type $(\mathcal{W},S)$.
    Let $D^{\prime}\in\Delta_{+}$ be a chamber in the positive half of the twin building which is stabilized by $g$, also let $D\in\Delta_{+}$ be an arbitrary chamber and $n\in\mathbb{Z}$. Then we know that
    $\delta_{+}(D^{\prime},D)=\delta_{+}(g^{n}.D^{\prime},g^{n}.D)=\delta_{+}(D^{\prime},g^{n}.D).$
    Next, denote with $l\colon\mathcal{W}\to\mathbb{N}$ the length function of the Weyl group $\mathcal{W}$ with respect to the generating set $S$. By [AB08, Definition 5.1 (WD3)] one derives a triangle inequality, which is given for the length function $l$, so that one can compute
    $l(\delta_{+}(D,g^{n}.D))\leq l(\delta_{+}(D,D^{\prime}))+l(\delta_{+}(D^{\prime},g^{n}.D))=2l(\delta_{+}(D,D^{\prime})).$
    Hence the orbit $\{g^{n}.D\mid n\in\mathbb{Z}\}$ is bounded and by the symmetry arguments of the to halves of a twin building, the claim follows.
[…]
    This statement is exactly the Bruhat-Tits fixed point theorem applied on the CAT(0) realization of $\Delta_{\pm}$, see [AB08, Corollary 12.67].
[…]
    As we have already seen, if $g$ stabilizes a chamber $D$, its inverse $g^{-1}$ also stabilizes $D$. The same calculation can be done to see, that if $g^{-1}$ stabilizes $D$, then also $g$ stabilizes $D$:
    $D=(gg^{-1}).D=g.(g^{-1}.D)=g.D.$
    Now let $R$ be a spherical residue stabilized by $g$ and by the computation
    $\Theta(R)=\Theta(g.R)=g^{-1}.\Theta(R),$
    and the previous comment, follows that $\Theta(R)$ is also stabilized by $g$. Hence the symmetric element $g$ stabilizes the spherical residue in each half of $\Delta_{\pm}$.
    The idea is now, to show that $g$ fixes a chamber in $R$ and hence a twin apartment in $\Delta_{\pm}$. In particular, if $g$ fixes a chamber $D$ in $R$, $g$ fixes also $\Theta(D)$, which is an opposite chamber. This leads to the fact, that $g$ fixes then the twin apartment $\Sigma(D,\Theta(D))$ point-wise, i.e. $g$ stabilizes a twin apartment chamberwise.
    By [AB08, Proposition 67.27], one knows that the stabilizer of a spherical residue of type $J$ equals a spherical parabolic subgroup $P_{J}$, compare this with the discussion before [Mar18, Proposition 7.75]. Since $g$ stabilizes $R$ and an opposite residue $\Theta(R)$, it is an element of
    $g\in P_{J}\cap\Theta(P_{J}).$
    Using the fact, that a parabolic subgroup, in the context of a Kac–Moody group, has a Levi decomposition (see [Rém02, 6.2.2 (Théorème)] or [Mar18, Proposition 7.75]) one can analyze this intersection. Therefore, denote by
    $P_{J}=G_{J}\ltimes U_{J}^{+}\,\text{ and }\Theta(P_{J})=G_{J}\ltimes U_{J}^{-}$
    the Levi decomposition of $P_{J}$ and note that $\Theta(P_{J})$ is obtained by applying the Chevalley involution.
    Here we used the notion from [Mar18, Chapter 7], in particular the group $G_{J}$ stands for a subgroup of $G$, which is generated by the torus $T$ and the root groups $U_{\alpha}$, where $\alpha$ are the roots contained in the span of simple roots corresponding to $J$, see the discussion before [Mar18, Proposition 7.75]. The second part of the Levi decomposition, $U_{J}^{\pm}$, is the normal subgroup of $U_{\pm}$ generated by $U_{\alpha}$, where $\alpha$ is in the complement of the roots used to generate $G_{J}$.
    Since $g$ is an element of the intersection, it follows that it can be written in two ways:
    $xu_{+}=g=yu_{-},$
    where $x,y\in G_{J}$, $u_{+}\in U_{J}^{+}$ and $u_{-}\in U_{J}^{-}$. In fact, one realizes that
    $g^{-1}.R=(u_{-}^{-1}y^{-1}).R=u_{-}^{-1}.R=R,$
    since $y^{-1}$ is an element of the stabilizer. It yields that $u_{-}\in P_{J}\cap U_{J}^{-}=\{e\}$, and analogously one can show with $\Theta(P_{J})$ that $u_{+}$ must also be trivial.
    And so, the action of $g$ on the residue depends only on the part of $G_{J}$, and since $\mathcal{W}_{J}$ is finite, this is a Chevalley group, i.e. a linear algebraic group. By [Bor91, Proposition 1.10] we can consider this group as a closed subgroup of $\operatorname{GL}_{n}(\mathbb{K})$. To clarify the notation, we write $M_{g}$ for the corresponding matrix in the subgroup of $\operatorname{GL}_{n}(\mathbb{K})$.
    The Chevalley involution translates for these matrices to transpose-inverse, composed with complex conjugation if $\mathbb{K}=\mathbb{C}$. Taking a closer look on the property that $g$ is symmetric turns $M_{g}$ it into a Hermitian matrix, in particular
    $\Theta(M_{g})=M_{g}^{-1}\Longleftrightarrow\left(\overline{M_{g}}\right)^{-T}=M_{g}^{-1}\Longleftrightarrow M_{g}^{T}=\overline{M_{g}}.$
    Thus, $M_{g}$ is diagonalizable and hence it is conjugated to an diagonal matrix $D_{g}$. Translated back into the language of Kac–Moody groups, this leads to the fact that $g$ is conjugate to an element of the standard torus $T$ of the Kac–Moody group $G$. Since the standard torus fixes the pair of fundamental chambers $(C_{+},C_{-})$ in $\Delta_{\pm}$, a conjugated element fixes a translated pair of the fundamental chambers. In detail if $T$ fixes $C_{\pm}$ and $aga^{-1}\in T$ denotes a conjugated element in $T$, one calculates
    $T.C_{\pm}=C_{\pm}\Longrightarrow\left(aga^{-1}\right).C_{\pm}=C_{\pm}\Longrightarrow g.\left(a^{-1}.C_{\pm}\right)=a^{-1}.C_{\pm}.$
    By the argument from above, $g$ fixes then the twin apartment $\Sigma(a^{-1}.C_{\pm},a^{-1}.\Theta(C_{\pm}))$ chamberwise. Note that all of these arguments apply to both halves of the twin building for reasons of symmetry.
The following lemma is a consequence of the latter statement and allows us to prove that a special quotient of a Kac–Moody group is a symmetric space. Thus, the lemma is one of the central parts of the theory we develop in this article.
The version for non-affine centered Kac–Moody groups over $\mathbb{R}$ can be found at [FHHK20, Proposition 3.38].
Lemma 4.3
Let $G$ be a split Kac–Moody group over $\mathbb{K}$ of type $\mathbb{A}$. Then
$\tau(G)\cap K=\{e\}.$
Proof:
Let $G$ be a Kac–Moody group and denote by $\Delta_{\pm}$ the associated twin building. Further, let $g$ be an element of $\tau(G)$, i.e. $g=h\Theta(h)^{-1}$, for an element $h\in G$. Then one calculates
$\Theta(g)=\Theta(h\Theta(h)^{-1})=\Theta(h)h^{-1}=g^{-1}.$
Thus, if $g\in K\cap\tau(G)$, one realizes that $g=\Theta(g)=g^{-1}$, from which it follows that $g$ is symmetric and has order at most $2$. This leads to the fact that the orbits of $g$ are bounded as in point 4 of 4.1. Hence by the proof of 1 from 4.1 one knows, that $g$ fixes the $\Theta$-stable twin apartment $\Sigma(C,\Theta(C))$ chamberwise. Thus $g$ is contained in a corresponding torus $T^{\prime}$ of $G$. By 3.15 it follows, that $K$ acts transitively on both halves of the associated twin building $\Delta_{\pm}$.
Hence there is a $k\in K$ such that $kT^{\prime}k^{-1}=T$. It follows that the element
$kgk^{-1}=k\tau(h)\Theta(k^{-1})=\tau(kh)$
is an element of $T$, $K$ and $\tau(G)$. By 3.14 and 3.21 follows that
$kgk^{-1}\in T\cap K\cap\tau(G)=\{e\}$
and thus $g=e$. ∎
4.2 Kac–Moody Symmetric Space
Loos’ abstract definition of a symmetric space, see 2.1, requires a set and a multiplication map $\mu$ that encodes point reflections within the symmetric space. It turns out that the Chevalley involution composed with the complex conjugation $\Theta$ (cf. Section 3.5) is essential to define such a map $\mu$. Since the Kac–Moody group $G$ is a topological group, see Section 3.2, we deduce that $K$ is a closed subgroup, since $\Theta$ is continuous and thus the quotient $G/K$ is a topological Hausdorff space.
Proposition 4.4
Let $G$ be a split Kac–Moody group over $\mathbb{K}$ of type $\mathbb{A}$. Define the reflection map
$\begin{split}\mu\colon G/K\times G/K&\to G/K\\ (gK,hK)&\mapsto\tau(g)\Theta(h)K.\end{split}$ (5)
The map $\mu$ is continuous, and if $\tau(G)\cap K=\{e\}$, then the pair $(G/K,\mu)$ is a symmetric space and there is a natural action by automorphisms
$G\to\operatorname{Sym}(G/K)\,,\;g\mapsto(hK\mapsto ghK).$
Proof:
The proof is a straightforward calculation, independent of the choice of $\mathbb{A}$ and $\mathbb{K}$, and can be found in [FHHK20, Proposition 4.2] or [Voc23, Proposition 4.3]. ∎
Definition 4.5
The symmetric space $(G/K,\mu)$ is the Kac–Moody symmetric space over $\mathbb{K}$ of type $\mathbb{A}$.
Remark 4.6
Note that this definition of the reflection map $\mu$ also holds for any topological group with an involution. Moreover, the first three properties in 2.1 can be proved without any assumption about the group, see the calculation in the proof of [FHHK20, Proposition 4.2]. Thus any topological group with an involution that satisfies the last property gives rise to a symmetric space.
  1. 1
    The first example which comes to mind is the group given by $(\mathbb{R}^{n},+)$ with involution $\Theta\colon\mathbb{R}^{n}\to\mathbb{R}^{n}$, $\Theta(x)=-x$. Then one computes for $\tau(x)=x-\Theta(x)=2x$ and hence the reflection of the symmetric space is given by $\mu(x,y)=\tau(x)+\Theta(y)+K=2x-y$ and $K$ is equal to $\{0\}\subseteq\mathbb{R}^{n}$.
  2. 2
    Another example is the Lie group $\operatorname{SL}_{n}(\mathbb{R})$ with involution $\Theta(A)=A^{-T}$. Clearly the fixed point set $K$ equals the special orthogonal group $\operatorname{SO}(n)$. Note, that $\tau(A)=AA^{T}$ for a matrix $A\in\operatorname{SL}_{n}(\mathbb{R})$, hence the matrices in the image of the twist map are symmetric and thus diagonalizable over $\mathbb{R}$. Moreover, the image $\tau(G)$ consists of positive semi-definite matrices.
    To check if $\tau(G)\cap K$ is trivial, we use that $\tau(A)$ can be diagonalized: let $S$ be a orthogonal real matrix such that
    $S(AA^{T})S^{-1}=D$
    is a diagonal matrix. Assume now, that $AA^{T}\in K=\operatorname{SO}(n)$, then it is true that
    $(AA^{T})(AA^{T})^{T}=\mathbb{E}_{n},$
    where $\mathbb{E}_{n}$ stands for the $n\times n$-identity matrix. Thus we can compute
    $\left(S(AA^{T})S^{-1}\right)^{2}$ $=D^{2}$
    $S(AA^{T})^{2}S^{-1}$ $=D^{2}$
    $SS^{-1}$ $=D^{2}$
    and it follows that the entries of $D$ are ±$\pm 1$. But by the positive semi-definite property only $+1$ is possible.
4.3 The Group Model
So far, we can describe a symmetric space associated with a Kac-Moody group by $(G/K,\mu)$, which is called the coset model. In this section we want to set up another model to describe the same symmetric space with $\tau(G)$ using the quadratic representation of an abstract symmetric space, see 2.4. The following section is based on [FHHK20, Section 4.12].
In order to use the quadratic representation, we need a pointed symmetric space. Therefore choose $eK$ as the base point. Then, by 2.4, there is an isomorphism of symmetric spaces
$(G/K,\mu)\cong(T(G/K,eK),\mu_{T}).$
Next, we take a closer look at $T(G/K,eK)$. Therefore, we first compute the kernel of the group action of $G$ on $G/K$, which is given by left multiplication:
$G\times G/K\to G/K\,,\;(g,hK)\mapsto ghK.$
If $g\in G$ is an element of the kernel, then $ghK=hK$ for all $h\in G$, in particular for $h\in K$ follows $ghK=K$ which implies that $g\in K$. Now assume that $g\in K$ acts trivially on $hK$ for an element $h\in G$. Then we can compute $ghK=hK\Leftrightarrow h^{-1}gh\in K$ and by using 3.14 (A1) it follows that $\tau(h^{-1}gh)=e$. By a direct calculation this leads to $g\tau(h)g^{-1}=\tau(h)$, and thus the kernel of the action is given by $C_{K}(\tau(G))$.
Next, denote by $\ell_{g}\colon G/K\to G/K$ the left multiplication with an element $g\in G$, i.e. $\ell_{g}(hK)=ghK$. It follows from the latter calculation that $\ell_{g}=\ell_{h}$ if $g$ and $h$ are in the same coset of $C_{K}(\tau(G))$. Then we can calculate
$s_{gK}(hK)$ $=\tau(g)\Theta(h)K$
$=\ell_{g}\left(\Theta(g^{-1}h)K\right)$
$=\ell_{g}\left(s_{eK}\left(g^{-1}hK\right)\right)$
$=\left(\ell_{g}\circ s_{eK}\circ\ell_{g^{-1}}\right)(hK).$
Using this, we can rewrite a transvection:
$s_{gK}\circ s_{eK}$ $=\ell_{g}\circ s_{eK}\circ\ell_{g^{-1}}\circ s_{eK}=\ell_{\tau(g)}$
and thus
$T(G/K,eK)$ =$=\{s_{gK}\circ s_{eK}\mid gK\in G/K\}$
=$=\{\ell_{\tau(g)}\mid g\in G\}$
=$=\{\tau(g)C_{K}(\tau(G))\mid g\in G\}.$
This leads to the observation, that the map
$\zeta\colon\tau(G)\to T(G/K,eK)$
is surjective. Furthermore, if an element is contained in the kernel, it must be contained in $K$ by the previous calculation. Since $\tau(G)\cap K=\{e\}$, see 4.3, it follows that the kernel of $\zeta$ is trivial. By transport of structure, the map establishes a symmetric space structure on $\tau(G)$, where the reflection map is given by $\mu_{\tau}(g,h)=gh^{-1}g$.
Altogether there are the following isomorphism of symmetric spaces
[…]
The symmetric space $(\tau(G),\mu_{\tau})$ is called group model of the Kac–Moody symmetric space. The reflection map $\mu_{\tau}$ is the same as the reflection map given in 2.2 4 for a general group.
Remark 4.7
As in [FHHK20, Proposition 4.11 (iii)], the transvection group of the Kac–Moody symmetric space can also be written as a quotient of the Kac–Moody group. But in the general case of arbitrary split Kac–Moody groups, we have to deal with the enlarged torus. We adopt the strategy used in the proof of [FHHK20, Proposition 4.11], but we have to pay attention to the torus.
Define the following quotient, which, as we will see, acts faithfully on a Kac–Moody symmetric space:
$\left(G^{\dagger}A\right)_{\operatorname{eff}}\coloneqq\nicefrac{{G^{\dagger}A}}{{C_{K}(G^{\dagger}A)}}.$
Since $M\subseteq K$ and $\tau(T)=A$, see 3.14, one concludes that
$\langle\tau(G)\rangle=\langle\tau(G^{\dagger}T)\rangle\subseteq G^{\dagger}A.$
On the other hand, one knows that $G^{\dagger}=\langle\tau(G^{\dagger})\rangle$ (cp. [FHHK20, Proposition 3.39]) as well as $\tau(T)=A$. Note that [FHHK20, Proposition 3.39] is formulated for centered Kac–Moody groups with an RGD system, and $G^{\dagger}$ is a centered Kac–Moody group with an RGD system, see the discussion after 3.13. Thus
$\langle\tau(G)\rangle=G^{\dagger}A.$
It follows from the previous observations that an element $g\in G$ acts trivially on a Kac–Moody symmetric space $G/K$ if and only if it centralizes $G^{\dagger}A$. In terms of the notation of left multiplication used above, $l(g_{1})=l(g_{2})$ only if $g_{1}$ and $g_{2}$ are in the same coset $C_{K}(G^{\dagger}A)$.
We now want to associate $\left(G^{\dagger}A\right)_{\operatorname{eff}}$ with the transvection group of $G/K$,
$\operatorname{Trans}(G/K)\coloneqq\langle s_{gK}\circ s_{hK}\mid gK,hK\in G/K\rangle.$
The strategy is the same as in the previous discussion and based on the proof of [FHHK20, Proposition 4.11]. Let $g\in G$ be an arbitrary element, $s_{gK}$ the corresponding point reflection and denote by $b=eK$ the base point of $G/K$. Then as done before, one calculates
$\begin{split}s_{gK}\circ s_{b}=l(\tau(g)).\end{split}$
Thus, for the generators of the transvection group, we observe the following presentation
$s_{gK}\circ s_{hK}$ $=\left(s_{gK}\circ s_{b}\right)\circ\left(s_{hK}\circ s_{b}\right)^{-1}$
$=l(\tau(g)\tau(h)^{-1}).$
Further, $\tau(h)^{-1}=\tau(\Theta(h))$, hence the transvection group is generated by elements of the form
$\tau(G)^{2}C_{K}\left(G^{\dagger}A\right).$
Together with 4.7 follows that
$\left(G^{\dagger}A\right)_{\operatorname{eff}}=\operatorname{Trans}(\nicefrac{{G^{\dagger}A}}{{K}})=\operatorname{Trans}(G/K).$
On the basis of this section, we have two isomorphic perspectives when we consider a symmetric Kac–Moody space. One is the coset model $G/K$, which is the natural way to construct the symmetric space according to Loos. The other is the group model $\tau(G)$, which allows us to view the symmetric space as a subset of the original Kac–Moody group $G$. This gives us an intrinsic, algebraic view of the symmetric space.
4.4 Flats
To analyze the structure of the Kac–Moody symmetric space of a Kac–Moody group $G$ over $\mathbb{K}$ of type $\mathbb{A}$, we first have to define flats. Euclidean flats are subsets of the symmetric space, which are isomorphic to the Euclidean symmetric space (cf. 2.2 1). In order to define Euclidean flats in the context of Kac–Moody symmetric spaces, we need to define a real form of the Cartan subalgebra $\mathfrak{h}=\sum_{1\leq i\leq n}\mathbb{C}\alpha^{\vee}_{i}+\sum_{n+1\leq i\leq 2n-l}\mathbb{C}h_{i}\cong\mathbb{C}^{2n-l}$:
$\mathfrak{a}\coloneqq\operatorname{span}_{\mathbb{R}}\left(\alpha^{\vee}_{1},\ldots,\alpha^{\vee}_{n},h_{n+1},\ldots,h_{2n-l}\right)\cong\mathbb{R}^{2n-l}.$
Proposition 4.8
Let $G$ be a split Kac–Moody group over $\mathbb{K}$ of type $\mathbb{A}$ and equip the real form $\mathfrak{a}$ with its Euclidean symmetric space structure (cp. 2.2 1). Then the map
$\begin{split}\varphi_{g}\colon\mathfrak{a}&\to gAK\\ X&\mapsto g\exp(X)K\end{split}$ (6)
is an isomorphism of symmetric spaces for every $g\in G$. In particular, the subset $gAK$ of the Kac–Moody symmetric space $G/K$ is an Euclidean flat of dimension $\dim\mathfrak{a}=2n-l$.
Proof:
First of all, it should be noted that by restricting the exponential mapping Section 3.2 to the real form $\mathfrak{a}$, $\exp_{\operatorname{KM}}(\mathfrak{a})$ is bijective and the image is given by the subgroup $A$ of the torus. Furthermore, in this case $\exp_{\operatorname{KM}}$ corresponds to the natural exponential map and in particular to the group homomorphism between $(\mathbb{R}^{2n-l},+)$ and $((\mathbb{R}_{>0})^{2n-l},\cdot)$.
Now, using that the action of $G$ on the symmetric space $G/K$ is given by an automorphism (cf. 4.4), we can assume that $g=e$. Let $X,Y\in\mathfrak{a}$ and use that for all $t\in A$ follows that $\Theta(t)=t^{-1}$ (cp. Section 3.5). Then one computes
$\mu(\varphi_{e}(X),\varphi_{e}(Y))$ $=\tau(\exp(X))\Theta(\exp(Y))K$
$=\exp(X)\Theta(\exp(X))^{-1}\Theta(\exp(Y))K$
$=\exp(2X-Y)K=\varphi_{e}(\mu_{Eucl}(X,Y)).$
This shows, that $\varphi_{g}$ is an isomorphism of symmetric spaces.
The last step is to check that $gAK\subset G/K$ is closed. 3.20 tells that the multiplication $m\colon U_{\pm}\times A\times K\to G$ is a homeomorphism. This yields that the subspace $AK$ and $gAK$, for any $g\in G$, are closed in $G/K$ since the multiplication with an element induces a homeomorphism within the topological group. ∎
Definition 4.9
Call the map $\varphi_{g}$ from the previous proposition chart centered at $g$ and the image $\varphi_{g}(\mathfrak{a})=gAK$ standard flat for every $g\in G$ of the Kac–Moody symmetric space $(G/K,\mu)$.
We conclude the section about flats by state a characterization. The following statement is based on [FHHK20, Theorem 5.17].
Theorem 4.10
Let $(G/K,\mu)$ be the Kac–Moody symmetric space with respect to a Kac–Moody group over $\mathbb{K}$ of type $\mathbb{A}$, where $\mathbb{A}$ is a symmetrizable, indecomposable generalized Cartan matrix. Then every weak flat is contained in a standard flat. Especially
  • standard flats are exactly the maximal flats;
  • all weak flats are Euclidean;
  • all weak flats are flats;
  • $G$ acts transitively on maximal flats.
Proof:
The proof of [FHHK20, Theorem 5.17] relies on the fact that a Kac–Moody group has an RGD system and that the coset model is isomorphic to the group model. The isomorphism between the group model and coset model is given in Section 4.3. The arguments in the proof of [FHHK20, Theorem 5.17] based on RGD systems are independent of the type of the generalized Cartan matrix $\mathbb{A}$ and of whether the underlying field $\mathbb{K}$ of the Kac–Moody group is $\mathbb{R}$ or $\mathbb{C}$. Therefore, the proof given for [FHHK20, Theorem 5.17] is applicable, with the exception of claim $5$.
In claim $5$ a reductive split real Lie group is used and the fact that it is a subgroup of $\operatorname{GL}_{m}(\mathbb{R})$. For $\mathbb{K}=\mathbb{C}$ we also have an algebraic group which is a subgroup of $\operatorname{GL}_{m}(\mathbb{C})$ (cp. [Bor91, Proposition 1.10]. As we have seen in the proof of 4.1, the involution $\Theta$ at the level of the matrix group can be understood as transpose, inversion, which is combined with complex conjugation (cp. [PR94, Theorem 3.7 and Proposition 3.9]). In the concrete calculation of the product $xy\cdot xy$, the upper left entry is equal to a sum of the squares of the absolute values of the entries in the first column. Now the arguments in claim $5$ work verbatim exactly as in the real case. ∎
4.4.1 Flats in the Group Model
This section is based on [FHHK20, Section 5.1].
Recall from Section 4.3 the isomorphism $\hat{\tau}\colon G/K\to\tau(G)$ between the coset and the group model and that it is equivariant under the action of $G$. We are now interested in the image of a standard flat in the group model under this isomorphism:
$\hat{\tau}(gAK)$ $=\tau(gA)$
$=g\tau(A)\Theta(g)^{-1}$
$=g\star\tau(A).$
Thus, the image of the standard flat $gAK$ in the group model is given by $g\star\tau(A)$. Therefore, denote by $F[g]\coloneqq g\star A\subseteq\tau(G)$ the standard flat of the group model (cp. [FHHK20, Proposition 5.4]). In order to establish a relationship between the flats in the symmetric space $G/K$, the twin apartments in the corresponding building $\Delta_{\pm}$ and the maximal tori of $G$, we need the following lemma.
Lemma 4.11 (Remark 5.5 from [FHHK20])
Let $G$ ba a split Kac–Moody group over $\mathbb{K}$ of arbitrary type. Then
$N_{K}(A)=N_{K}(T).$
Proof:
Note that $N_{K}(T)<N_{K}(A)$, based on the fact that the conjugation is continuous and $A$ is the identity component of $T$.
For the reverse inclusion, consider an element normalizing $A$, then the conjugation with this element shifts $T$ to a maximal torus $T^{\prime}$ of $G$. But the standard maximal torus $T$ containing $A$ is unique ([Cap09, Lemma 4.9]). It follows that $T=T^{\prime}$, thus this element normalizes $T$. ∎
Proposition 4.12 (Remark 5.5 from [FHHK20])
There is a $G$-equivariant bijection between the set of maximal tori of $G$, the set of standard flats in $G/K$ and the twin apartments of $\Delta_{\pm}$, the corresponding twin building that belongs to $G$.
In fact, $G$ acts transitively on each of these objects.
Proof:
First denote by $T_{max}$ the set of all maximal tori of $G$ and with $F_{std}$ the set of all standard flats in $G/K$. Then there is a $G$-equivariant map between these sets given by
$f\colon T_{max}$ $\to F_{std}$
$gTg^{-1}$ $\mapsto gAK.$
To see that this map is well-defined, we first look at $N_{G}(T)$. We know by [FHHK20, Lemma 3.27] that $N_{G}(T)=A\ltimes N_{K}(T)$, since the proof of the quoted statement is given by 3.14 and the refined Birkhoff decomposition (see [Kum02, Theorem 5.2.3 (g)] or proof of 3.21). Together with 4.11 and [FHHK20, Lemma 3.27] follows that $N_{G}(T)$ stabilizes $AK$: Let $g\in N_{G}(T)$, then it can be written as $g=ak$ for $a\in A$ and $k\in N_{K}(A)$ and hence one calculates
$gAK=akAK=a(kA^{-1}k)kK=aAK=AK.$
Together with the fact that $N_{G}(T)$ acts trivially by conjugation on $T_{max}$, it follows that $G/N_{G}(T)$ can be identified with $T_{max}$ by mapping $gN_{G}(T)\mapsto gTg^{-1}$, in particular $f$ is surjective.
For the injectivity, we use the group model and compute
$F[g]\Theta(F[g])$ $=g\tau(A)\Theta(g)^{-1}\Theta(g)\Theta(\tau(A))g^{-1}$
$=g\tau(A)\Theta(\tau(A))g^{-1}$
$=g\tau(A)g^{-1}=gAg^{-1}.$
Since $A$ is contained in the standard maximal torus $T$ (see [Cap09, Lemma 4.9]), $gAg^{-1}$ is contained in $gTg^{-1}$.
Thus $gTg^{-1}$ is uniquely determined by the flat $F[g]$, and the map $f$ is injective.
Note that $T$ is the stabilizer of the standard twin apartment and the group $G$ acts transitively on $T$ and the twin apartments. For the action of $G$ on $T$ the stabilizer of the standard torus is given by $N_{G}(T)$ and for the action of $G$ on the twin apartments is $N_{G}(T)$ the stabilizer of the standard twin apartment. Thus, for the twin apartments follows that the maximal tori of $G$ are exactly the chamberwise stabilizers of the twin apartments of $\Delta_{\pm}$. ∎
5 Automorphism of Kac–Moody Symmetric Spaces
As a next step in developing the theory of symmetric spaces for Kac–Moody groups over $\mathbb{K}$ of type $\mathbb{A}$, one can analyzes the automorphism group of the whole symmetric space and the local action on flats. For the non-affine real case, in chapter 6 of [FHHK20], the authors observe several statements for the global automorphism group on the symmetric space as well as for the local automorphism group with respect to the flats.
5.1 Global Action
From the characterization of flats one can deduce that the Kac–Moody group acts transitively on pointed maximal flats. In the following we call such an action of a Kac–Moody group $G$ on the corresponding symmetric space strongly transitive, see [FHHK20, Definition 2.32].
Proposition 5.1
Let $G$ be Kac–Moody group over $\mathbb{K}$ of arbitrary type. Then $G$ acts strongly transitively on the symmetric space $G/K$.
Proof:
We are going to use [FHHK20, Proposition 2.33]: If a subgroup of the automorphism group of the symmetric space acts, which has to contain the transvection group, transitively on the maximal flats, then the group acts strongly transitively. This statement is shown for abstract symmetric spaces, hence we can apply it here. In 4.7 we have seen that $G$ contains the transvection group, so we can use the statement quoted.
By 4.12 we already know that $G$ acts transitively on standard flats and by 4.10 all standard flats are maximal flats. Therefore, the statement follows. ∎
A central assumption in [FHHK20] which allows to decompose an automorphism, is that the Kac–Moody group is centered (cp. 3.13), i.e. the Kac–Moody group is generated only by its root groups. To establish a similar results for non-centered Kac–Moody groups $G$ of type $\mathbb{A}$, we have to take care of the extended torus. In order to do this, we consider in the following section $G^{\dagger}\trianglelefteq G$.
By 3.13 we know, that the subgroup $G^{\dagger}$ is a centered Kac–Moody group, thus using [CM05, Theorem 2.4], we can write any automorphism of $G^{\dagger}$ as a product of an inner automorphism, the Chevalley involution, a diagonal automorphism, a graph automorphism and a field automorphism. Following, we give a brief overview of these automorphism, for details see [CM05, Section 8.2].
To do this, we need to recall some thins from Section 3.2. First, the map $\eta$ from the Tits functor, further the fact, that an element of the root group $U_{\alpha}$ is given by $x_{\alpha}(r)$ for $r\in\mathbb{K}\in\{\mathbb{R},\mathbb{C}\}$ and $\alpha\in\Delta^{re}$. And lastly, that the torus of a Kac–Moody group $T=\mathfrak{T}_{\Lambda}(\mathbb{K})$ is isomorphic to $\operatorname{Hom}_{grp}(\Lambda,\mathbb{K}^{\times})$. Hence we can define the map
$\iota(t)\colon\Lambda\to\mathbb{K}^{\times}\,,\;\lambda\mapsto t(-\lambda)=(t(\lambda))^{-1}$
for each $t\in T$. Now, for a given centered split Kac–Moody group $H$ of type $\mathbb{A}$, $r\in\mathbb{K}$, $\alpha_{i}\in\Pi$, $\alpha\in\Delta^{re}$ and $t\in T$, one can describe the types of the automorphisms as follows:
  1. 1
    An inner automorphism $c_{g}\colon H\to H$ for an element $g\in H$ is given by $c_{g}(h)=ghg^{-1}$.
  2. 2
    A sign automorphism is a map $s\colon H\to H$ such that
    $s(x_{\alpha_{i}}(r))=x_{-\alpha_{i}}(r)\,,\;s(x_{-\alpha_{i}}(r))=x_{\alpha_{i}}(r)\,\text{ and }\;s(\eta(t))=\eta(\iota(t)).$
  3. 3
    A diagonal automorphism is the identity on the torus and acts on $x_{\alpha}(r)$ by multiplying $r$ with an element of the field depending on $\alpha\in\Delta^{re}$.
  4. 4
    Let $\sigma$ be a permutation of the index set $I$, then a diagram automorphism is given as a map $\gamma$ such that
    $\gamma(x_{\alpha_{i}}(r))=x_{\alpha_{\sigma(i)}}(r)\,,\;\gamma(x_{-\alpha_{i}}(r))=x_{-\alpha_{\sigma(i)}}(r)\,\text{ and }\;\gamma(\eta(r^{h_{i}}))=\eta(r^{h_{\sigma(i)}}).$
  5. 5
    A field automorphism $f\colon\mathbb{K}\to\mathbb{K}$ acts on $H$ using the fact that $H$ is defined by a functor. In particular for any $t\in T$ it is $f\circ t$ and for any $\alpha\in\Delta^{re}$ and $r\in\mathbb{K}$ it is $f(x_{\alpha}(r))=x_{\alpha}(f(r))$.
Note that we choose $H$ as $G^{\dagger}$ for our purpose.
Remark 5.2
By the discussion in [CM05, Section 8.2] it follows that the sign automorphism is unique and from [Cap09, Section 8.2.1] follows that it is related to the Chevalley involution. It follows that for $\mathbb{K}=\mathbb{R}$ one has
$s=\Theta$
and for $\mathbb{K}=\mathbb{C}$ one has
$s\circ f=\Theta,$
where $f$ denotes the field automorphism of the complex conjugation.
With this concrete decomposition, we can now rewrite [FHHK20, Proposition 6.5] in the context of the centered subgroup $G^{\dagger}$ of a Kac–Moody group $G$ of type $\mathbb{A}$ over $\mathbb{R}$ or $\mathbb{C}$.
Proposition 5.3
Let $G$ be a split Kac–Moody group of type $\mathbb{A}$ over $\mathbb{K}\in\{\mathbb{R},\mathbb{C}\}$ and $G^{\dagger}\trianglelefteq G$ the centered subgroup. Then one can decompose the automorphism group, i.e.
$\operatorname{Aut}(G^{\dagger})\cong\left(\operatorname{Ad}(G^{\dagger})\rtimes\left(D\times\langle\Theta\rangle\right)\right)\rtimes\operatorname{Aut}(\Gamma),$
where $\operatorname{Ad}(G^{\dagger})\coloneqq G^{\dagger}/Z(G^{\dagger})$, $D$ is a group of diagonal automorphisms and $\operatorname{Aut}(\Gamma)$ denotes the diagram automorphisms of the corresponding Dynkin diagram with respect to $\mathbb{A}$.
Note that this explicit statement about the structure of the automorphism group is mainly based on the result of [Cap09, Theorem 2.4].
Proof:
The case $k=\mathbb{R}$ is proven in [FHHK20, Proposition 6.5], where it is formulated for any generalized Cartan-Matrix $\mathbb{A}$.
For the complex case, we recall the decomposition of an automorphism of $G^{\dagger}$ ([CM05, Theorem 2.4]: Denote with $c_{g}$ an inner automorphism, with $\omega$ the Chevalley involution, with $d$ a diagonal automorphism, with $\gamma$ a diagram automorphism and with $f$ a field automorphism. Then any automorphism $\alpha\in G^{\dagger}$ can be written as
$\alpha=c_{g}\circ\omega\circ d\circ\gamma\circ f.$
In the complex case, there is only one non-trivial continuous field automorphism $f$, the complex conjugation. In particular, $f$ commutes with all other automorphism in this composition and the generated subgroups of these homomorphisms intersect trivially. By the definition of the involution $\Theta$ follows $\omega\circ f=\Theta$. ∎
Remark 5.4
  • Since we are interested in statements about Kac–Moody symmetric spaces, we also need to define a symmetric space with respect to the centered subgroup $G^{\dagger}$. In particular, let $K^{\dagger}\coloneqq G^{\dagger}\cap K$, and since $\Theta$ is an automorphism of $G$, it follows that $(G^{\dagger})^{\Theta}=K^{\dagger}$. Now define the corresponding Kac–Moody symmetric space
    $\left(\nicefrac{{G^{\dagger}}}{{K^{\dagger}}},\mu\mid_{G^{\dagger}}\right)$
    and denote it by $X^{\dagger}$. Note that this is the unreduced Kac–Moody symmetric space, which is also treated in part in the work on which this is based, see [FHHK20, Definition 4.5 (i)].
  • Since $G^{\dagger}$ is a centered Kac–Moody group which corresponds to the Kac–Moody group treated in [FHHK20], we can apply [FHHK20, Proposition 6.4]. The proposition states that
    $\operatorname{Aut}(G^{\dagger})\hookrightarrow\operatorname{Aut}(\Delta)$
    which allows us to interpret automorphisms of $G^{\dagger}$ as automorphisms of the corresponding twin building $\Delta_{\pm}$ from the RGD system associated to the Kac–Moody group $G$. By an automorphism of $\Delta$ we mean here a map of the chamber set $\Delta=\Delta_{+}\cup\Delta_{-}$ into itself, preserving the adjacency relation and the opposition relation.
We conclude this section by pointing out the following statement, which describes the automorphism group of a Kac–Moody symmetric space coming from the centered subgroup. Since $X^{\dagger}$ corresponds to the unreduced Kac–Moody symmetric space in [FHHK20], the proof of this theorem is the same as that of the quoted theorem. Note that the underlying field plays no role in the proof of this statement.
Theorem 5.5 (Theorem 6.12 (i) [FHHK20])
Let $G$ be a split Kac–Moody group of type $\mathbb{A}$ over $\mathbb{K}\in\{\mathbb{R},\mathbb{C}\}$ and $G^{\dagger}\trianglelefteq G$ its centered subgroup. Then
$\operatorname{Aut}\left(X^{\dagger}\right)\cong\operatorname{Aut}_{\text{eff}}(G^{\dagger})\coloneqq\left(\nicefrac{{G^{\dagger}}}{{C_{K^{\dagger}}(G^{\dagger})}}\rtimes(D\times\langle\Theta\rangle)\right)\rtimes\operatorname{Aut}(\Gamma),$
where $C_{K^{\dagger}}(G^{\dagger})$ is the centralizer of $G^{\dagger}$ in $K^{\dagger}$.
6 Local Action
One can define local transformations of flats for general symmetric spaces (cf. [FHHK20, Section 2.31]) and then one can transfer these definitions into the context of Kac–Moody symmetric spaces and analyze them.
In this section we denote by $G$ a split Kac–Moody group over $\mathbb{K}\in\{\mathbb{R},\mathbb{C}\}$ of type $\mathbb{A}$, where $\mathbb{A}$ denotes an indecomposable, symmetrizable generalized Cartan matrix of arbitrary type. Further denote by B the bilinear form on the Cartan subalgebra $\mathfrak{h}$ induced by the symmetrizable generalized Cartan matrix. For details of the construction of this bilinear form, see [Kac90, §2.1] (or [Mar18, Section 3.7] or [FHHK20, A.22]).
This section is primarily based on [FHHK20, Section 6.14]. We will sometimes refer to statements from [FHHK20] without justifying why, since the statements we refer to do not take into account the underlying field or type of the Kac–Moody group. If there is a statement that does take this into account and needs further explanation before we apply it in our context, we will point it out.
For an abstract symmetric space according to Loos one can define the following terms, compare them with [FHHK20, Definition 2.32] and [FHHK20, Definition 2.35]: Let $X$ be an abstract symmetric space containing a maximal Euclidean flat $F$ of rank $m$ and denote by $p$ a point of $F$.
  • The pair $(p,F)$ is called a pointed maximal flat.
  • Call a point $q\in F$ singular with respect to $p$, if there exists a second maximal flat $E\subseteq X$ such that $p,q\in E$ and $E\neq F$. If not, call $q$ regular. Further, let ${F^{\text{sing}}}(p)$ denote the subset of singular points of $(p,F)$ and ${F^{\text{reg}}}(p)$ the set of regular points. It follows that the flat $F$ can be written as $F={F^{\text{reg}}}(p)\sqcup{F^{\text{sing}}}(p)$.
  • The isomorphism of symmetric spaces $\varphi\colon\mathbb{R}^{m}\to F$, see the discussion before [FHHK20, Proposition 2.21], is called a chart of $F$ centered at $p$ if $p=\varphi(0)$.
  • Call a map $f\colon F\to F$ linear at $p$ if for a chart of $F$ centered at $p$ the following is true
    $\varphi^{-1}\circ f\circ\varphi\in\operatorname{GL}_{m}(\mathbb{R}).$
  • A map $f\colon F\to F$ linear at $p$ is called local transformation of a pointed flat $(p,F)$ if it preserves the decomposition
    $F={F^{\text{reg}}}(p)\sqcup{F^{\text{sing}}}(p).$
    The set of all local transformations of $(p,F)$ forms a group and is denoted by $\operatorname{GL}(p,F,{F^{\text{sing}}})$.
The idea in the following is to use these terms in the context of Kac–Moody symmetric spaces, since we have a good understanding of flats there. In [FHHK20, Section 6.14] the authors give a description of singular points in terms of the real form $\mathfrak{a}$ and connect it to the notion of singular points of abstract symmetric spaces. Afterwards they analyze the group of local transformations and compare the Weyl group of the Kac–Moody symmetric space with the Weyl group of the Kac–Moody algebra.
The difficulty in applying this description to the general case for all generalized Cartan matrices $\mathbb{A}$ is that in [FHHK20] the semisimple adjoint quotient of a Kac–Moody group is used (cf. [FHHK20, Definition 3.8]) and for Kac–Moody groups of affine type the quotient prevents a representation of the Weyl group from being faithful and the corresponding bilinear form on $\mathfrak{h}$ from being non-degenerate (cf. [FHHK20, Section A.22]). In the following we give a general definition of the representation which is faithful for real or complex split Kac–Moody groups of arbitrary type.
Recall that the Weyl group $\mathcal{W}$ of the Kac–Moody algebra $\mathfrak{g}(\mathbb{A})$ is a subgroup of $\operatorname{GL}(\mathfrak{h})$. As the reflections depend on the generalized Cartan matrix $\mathbb{A}$, which has integer entries, the action stabilizes the real form $\mathfrak{a}$, so we can define the Kac–Moody representation
$\begin{split}\rho_{KM}\colon\mathcal{W}\to\operatorname{GL}(\mathfrak{a}).\end{split}$ (7)
According to [Kac90, Proposition 3.12], this representation is faithful and the invariant symmetric bilinear form B on $\mathfrak{h}$ is non-degenerate, see [Kac90, Lemma 2.1 b)] (or [Kac90, Theorem 2.2] or [FHHK20, Proposition A.23]). Note that the generating reflections $r_{i}^{\vee}\in\mathcal{W}$ can also be defined using the bilinear form B, see [FHHK20, A.22].
Since the elements of $\mathcal{W}$ act as linear reflections on $\mathfrak{a}$ under the representation $\rho_{KM}$, we can define a reflection hyperplane with respect to a real root $\alpha\in\Delta^{re}$ as
$H_{\alpha}\coloneqq\operatorname{Fix}(\rho_{KM}(r_{\alpha}^{\vee}))\subseteq\mathfrak{a}.$
From the non-degeneracy of B on $\mathfrak{h}$ it follows that B is also non-degenerate on $\mathfrak{a}$. Therefore, the reflection $r_{\alpha}^{\vee}$ under $\rho_{KM}$ is the only B-orthogonal reflection on the hyperplane $H_{\alpha}$. In particular, this leads to an one-to-one correspondence between the real roots $\alpha\in\Delta^{re}$ and the hyperplanes $H_{\alpha}$.
Using these concepts, we can define singular points in terms of the real form $\mathfrak{a}$ as the union of the reflection hyperplane:
${\mathfrak{a}^{\text{sing}}}\coloneqq\bigcup_{\alpha\in\Delta^{re}}H_{\alpha}.$
As we saw at the beginning of this section, we can also define singular points on $\mathfrak{a}$ by using the notion of singular points in the context of abstract symmetric spaces. To be able to compare the two terms, we need to make them compatible. Therefore, let $(e,AK)$ be the standard flat of the symmetric space $G/K$. Using the chart $\varphi_{e}$, one can move the singular points of the flat $AK$ into the real form $\mathfrak{a}$, i.e.
$\varphi_{e}^{-1}\left({(AK)^{\text{sing}}}(e)\right),$
where ${(AK)^{\text{sing}}}(e)$ denotes the set of singular points of $(e,AK)$.
Now we can state that the descriptions of singular points on the real form are identical. The statement for non-affine real Kac–Moody symmetric spaces can be found in [FHHK20, Proposition 6.16].
Proposition 6.1
Let $G/K$ be a Kac–Moody symmetric space. Then
$\bigcup_{\alpha\in\Delta^{re}}H_{\alpha}={\mathfrak{a}^{\text{sing}}}=\varphi_{e}^{-1}\left({(AK)^{\text{sing}}}(e)\right).$
Proof:
We can use the proof of [FHHK20, Proposition 6.16] verbatim, since it starts with an assertion that follows in our general case from 5.1, and then uses only building-theoretic arguments that are available since each Kac–Moody group has an RGD system over $\mathbb{K}\in\{\mathbb{R},\mathbb{C}\}$, see 3.10. Note that the type of $\mathbb{A}$ does not matter. ∎
We are using this statement to analyze the group of local transformations. Therefore, define the group
$\operatorname{GL}(\mathfrak{a},{\mathfrak{a}^{\text{sing}}})\coloneqq\{f\in\operatorname{GL}(\mathfrak{a})\mid f({\mathfrak{a}^{\text{sing}}})={\mathfrak{a}^{\text{sing}}}\},$
which preserves the decomposition of $\mathfrak{a}$ into singular and regular points. By 6.1 we can deduce that there is an isomorphism
$\operatorname{GL}(e,AK,{(AK)^{\text{sing}}})$ $\to\operatorname{GL}(\mathfrak{a},{\mathfrak{a}^{\text{sing}}})$
$f$ $\mapsto\varphi_{e}^{-1}\circ f\circ\varphi_{e}.$
Since $G$ acts strongly transitively, see 5.1, any pointed maximal flat $(p,F)$ can mapped to the standard pointed flat $(e,AK)$. Furthermore, the $G$ action preserves the decomposition into regular and singular points. Thus we observe an isomorphism
$\operatorname{GL}(p,F,{F^{\text{sing}}})$ $\to\operatorname{GL}(\mathfrak{a},{\mathfrak{a}^{\text{sing}}})$
$f$ $\mapsto\varphi_{p}^{-1}\circ f\circ\varphi_{p},$
where $\varphi_{p}$ stands for a chart centered at $p$ and by 6.1 the chart identifies ${\mathfrak{a}^{\text{sing}}}$ with ${F^{\text{sing}}}(p)$.
Convention 6.2
In order to be able to use some corollaries from the appendix of [FHHK20], we need an additional assumption in the rest of this section: the generalized Cartan matrix $\mathbb{A}\in\mathbb{Z}^{n\times n}$ must be symmetrizable and $n$ must be at least $2$.
By Corollary A.37 from [FHHK20] follows that we can write any element of $\operatorname{GL}(\mathfrak{a},{\mathfrak{a}^{\text{sing}}})$ as a product of a B-orthogonal linear transformation and a homothety.
Note that [FHHK20, Corollary A.37] is applicable here, since the arguments and statements they used in the proof of this corollary are based on the fact that the bilinear form B coming from the generalized Cartan matrix $\mathbb{A}$ is non-degenerate and that $\rho_{KM}$ is faithful. All these properties are also satisfied in our context and thus we can apply the corollary.
Hence, by the isomorphism between $\operatorname{GL}(p,F,{F^{\text{sing}}})$ and $\operatorname{GL}(\mathfrak{a},{\mathfrak{a}^{\text{sing}}})$, we can use this decomposition of $\operatorname{GL}(\mathfrak{a},{\mathfrak{a}^{\text{sing}}})$ to rewrite the group of local transformations.
Therefore given a pointed flat $(p,F)$ and a local transformation $f\in\operatorname{GL}(p,F,{F^{\text{sing}}})$, then $f$ is a local automorphism if $\varphi^{-1}\circ f\circ\varphi$ is B-orthogonal for a chart $\varphi$ centered at $p$. We denote the subgroup of local automorphism by $\operatorname{Aut}(p,F)<\operatorname{GL}(p,F,{F^{\text{sing}}})$. By the previous discussion, we can decompose any local transformation into a product of a homothety and a B-orthogonal transformation. It follows that we can now decompose the group of local transformation as follows
$\operatorname{GL}(p,F,{F^{\text{sing}}})\cong\mathbb{R}_{>0}\times\operatorname{Aut}(p,F),$ (8)
where $\mathbb{R}_{>0}$ represents the action of the homotheties on the flat $F$.
Proposition 6.3 (cf. Corollary 6.20 in [FHHK20])
Let $G$ be a Kac–Moody group over $\mathbb{K}$ of type $\mathbb{A}$, where $\mathbb{A}\in\mathbb{Z}^{n\times n}$ is non-spherical and symmetrizable. For any pointed flat $(p,F)$ one has
$\operatorname{Aut}(p,F)\cong(\mathcal{W}\rtimes\operatorname{Aut}(\mathcal{W},S))\times\mathbb{Z}/2\mathbb{Z}$
and hence
$\operatorname{GL}(p,F,{F^{\text{sing}}})\cong\mathbb{R}_{>0}\times(\mathcal{W}\rtimes\operatorname{Aut}(\mathcal{W},S))\times\mathbb{Z}/2\mathbb{Z},$
where $\operatorname{Aut}(\mathcal{W},S)$ denotes the group of automorphisms of $\mathcal{W}$ that preserve $S$ as a set.
Proof:
Recall that the Weyl group $\mathcal{W}$ of the Kac–Moody group with generating set $S$ is a Coxeter system $(\mathcal{W},S)$ and that a Coxeter system induces a simplicial Coxeter complex $\Sigma$ (cf. [AB08, Chapter 3] and [FHHK20, Section A.1]). Denote in the following with $\operatorname{Aut}(\Sigma)$ the group of simplicial automorphisms. By [FHHK20, Lemma A.31] one knows that
$\operatorname{Aut}(\Sigma)=W\rtimes\operatorname{Aut}(\mathcal{W},S)$
and there the action of $\operatorname{Aut}(\Sigma)$ on $\mathfrak{a}$ is also described.
Next denote by $\operatorname{O}(\mathfrak{a},\textbf{B})$ the group of orthogonal linear maps with respect to the bilinear form B and define the group
$\operatorname{O}(\mathfrak{a},{\mathfrak{a}^{\text{sing}}})\coloneqq\operatorname{O}(\mathfrak{a},\textbf{B})\cap\operatorname{GL}(\mathfrak{a},{\mathfrak{a}^{\text{sing}}}).$
By [FHHK20, Proposition A.32] and [FHHK20, Remark A.33 (ii)], one derives that
$\operatorname{Aut}(\Sigma)\times\mathbb{Z}/2\mathbb{Z}\cong\operatorname{O}(\mathfrak{a},{\mathfrak{a}^{\text{sing}}})$
and hence that
$\operatorname{Aut}(\Sigma)\times\mathbb{Z}/2\mathbb{Z}\cong\operatorname{Aut}(p,F).$
Note that the quoted statements apply to Kac–Moody groups of arbitrary type, since they require a faithful representation $\rho_{KM}$ and a non-degenerate bilinear form B on $\mathfrak{a}$. Both are given in the general context.
Together with [FHHK20, Lemma A.31], we obtain the first isomorphism. The second isomorphism follows directly from Equation 8. ∎
6.1 Comparing the Weyl Groups
As at the end of [FHHK20, Section 2], one can define a Weyl group for an abstract symmetric space. In this section we will relate this Weyl group of the symmetric space to the Weyl group of the Kac–Moody group. The section follows [FHHK20, Section 2.31].
Let $(X,\mu)$ be a symmetric space containing a maximal Euclidean flat, let $G<\operatorname{Aut}(X)$ be a group acting transitively on maximal flats of $X$, and let $(p,U)$ be a maximal pointed flat. Then define the following terms:
$\operatorname{Stab}_{G}(p,U)\coloneqq\{g\in G\mid g(U)=U\,,\;g(p)=p\}$
and
$\operatorname{Fix}_{G}(p,U)\coloneqq\{g\in G\mid\forall u\in U:g(u)=u\}.$
Then one can define the geometric Weyl group by
$W(G\curvearrowright X)\coloneqq\nicefrac{{\operatorname{Stab}_{G}(p,U)}}{{\operatorname{Fix}_{G}(p,U)}},$
see [FHHK20, Definition 2.37 (i)].
Note that by [FHHK20, Proposition 2.36 (i)] the definition of the geometric Weyl group is independent of the choice of the pointed flat up to conjugation.
Now we are able to compare the Weyl group of a Kac–Moody symmetric space with the Weyl group of the underlying Kac–Moody group. Therefore, denote in the following with $G$ a real Kac–Moody group of arbitrary type $\mathbb{A}$. The reason why we consider a real Kac–Moody group goes back to the structure of the subgroup $M$ of the torus $T$ (see 3.14). In the following discussion we will see that it is necessary that the elements of $M$ have order $2$, which is true for real Kac–Moody groups.
Recall from the discussion after 3.11 that the centered subgroup $G^{\dagger}\trianglelefteq G$ leads to a twin building of the same type as $G$ and hence that the Weyl groups are equal, i.e.
$\nicefrac{{N_{G^{\dagger}}\left(T^{\dagger}\right)}}{{T^{\dagger}}}\cong\mathcal{W}\cong\nicefrac{{N_{G}(T)}}{{T}}.$ (9)
Recall that the extended Weyl group $\widetilde{\mathcal{W}}\coloneqq\langle\tilde{s}_{i}\mid i\in I\rangle$ (cp. Section 3.2) surjects on $\mathcal{W}$.
Due to [KP85, (2.4)] (see also [Mar18, Exercise 7.49]), the elements $\tilde{s}_{i}^{2}=(-1)^{h_{i}}$ are contained in the torus and thus they generate a subgroup of $\widetilde{\mathcal{W}}\cap T$ which has order $2^{n}$. In addition, according to [KP85, Corollary 2.3 (a)] the elements of the intersection $\widetilde{\mathcal{W}}\cap T$ have an order less than or equal to $2$.
Note that the standard torus of $G^{\dagger}$ is denoted by $T^{\dagger}$ and can be decomposed as a topological group into
$T^{\dagger}=M^{\dagger}\times A^{\dagger},$
where $A^{\dagger}$ and $M^{\dagger}$ are defined as in 3.14. By the construction of $G^{\dagger}$, it follows that $T^{\dagger}\cong(\mathbb{R}^{\times})^{n}$, where $n$ is the size of the Cartan matrix $\mathbb{A}$. Hence it follows by the construction of $M^{\dagger}$ (cp. 3.14) that $M^{\dagger}$ is the unique maximal finite subgroup of elements of order $2$. This implies the following
$\widetilde{\mathcal{W}}\cap T=M^{\dagger}.$
Further, by the definition of $\widetilde{\mathcal{W}}$ it follows that the extended Weyl group is a subgroup of the centered subgroup $G^{\dagger}$. In Section 3.2 we have seen that $\widetilde{\mathcal{W}}$ surjects through $\operatorname{Ad}(-)$ restricted to $\mathfrak{h}$ onto $\mathcal{W}$. Moreover, the kernel of this surjective map is exactly the intersection of $T$ with $\widetilde{\mathcal{W}}$. Thus we obtain the isomorphism
$\nicefrac{{\widetilde{\mathcal{W}}}}{{(\widetilde{\mathcal{W}}\cap T)}}\cong\mathcal{W}.$
Alternatively, we can also use [KP85, Proposition 2.1] to obtain the same statement.
Altogether it follows that
$\nicefrac{{\widetilde{\mathcal{W}}}}{{M^{\dagger}}}\cong\mathcal{W}.$ (10)
To compare the Weyl groups we need the following helpful lemma, which is a generalization of [FHHK20, Corollary 3.28].
Lemma 6.4 (cf. Corollary 3.28 from [FHHK20])
Let $G$ be a split Kac–Moody group of arbitrary type $\mathbb{A}$. Then the extended Weyl group satisfies
$\widetilde{\mathcal{W}}M=N_{K}(T).$
Proof:
Consider the canonical projection
$\pi\colon N_{G}(T)\to\nicefrac{{N_{G}(T)}}{{T}},$
and recall from 3.14 that $M\subset K$. In the proof of [FHHK20, Lemma 3.27 (ii)], using the Iwasawa decomposition and the refined Birkhoff decomposition, it is shown that $N_{G}(T)=N_{K}(T)T$ holds. Note that the refined Birkhoff decomposition can be used since the $BN$-pair obtained from $G$ is refined according to [Rém02, Theorem 1.5.4]. The restriction $\pi\mid_{N_{K}(T)}$ is therefore still surjective. From [FHHK20, Lemma 3.27 (ii)] it now follows that
$\nicefrac{{N_{G}(T)}}{{T}}=\nicefrac{{A\rtimes N_{K}(T)}}{{(A\rtimes M)}}\cong\nicefrac{{N_{K}(T)}}{{M}},$
and therefore the kernel of $\pi\mid_{N_{K}(T)}$ is equal to $M$. Thus
$\nicefrac{{N_{K}(T)}}{{M}}\cong\nicefrac{{N_{G}(T)}}{{T}}\cong\mathcal{W}.$
On the other hand, one knows that
$\nicefrac{{\widetilde{\mathcal{W}}M}}{{M}}\cong\nicefrac{{\widetilde{\mathcal{W}}}}{{(\widetilde{\mathcal{W}}\cap M)}}=\nicefrac{{\widetilde{\mathcal{W}}}}{{M^{\dagger}}}\cong\mathcal{W}$
and thus
$\nicefrac{{\widetilde{\mathcal{W}}M}}{{M}}\cong\nicefrac{{N_{K}(T)}}{{M}},$
which shows the desired statement. ∎
Finally, we can give the statement which compares the Weyl group of a real Kac–Moody group with the Weyl group of the corresponding symmetric space.
Proposition 6.5
Let $G$ be a split Kac–Moody group over $\mathbb{R}$ and $(p,F)$ a pointed flat in the symmetric space $G/K$. Then $\operatorname{Stab}_{G}(p,F)\cong N_{K}(T)$ and $\operatorname{Fix}_{G}(p,F)\cong M$, moreover
$W\left(G\curvearrowright G/K\right)\cong\mathcal{W}.$
Proof:
Since the definition of the Weyl group of a symmetric space is independent of the choice of the pointed flat, we can choose the standard flat $(e,AK)$.
Now we analyze the stabilizer and fixator of $G/K$ with respect to the action of $G$ on $G/K$. Observe that the fixator of $e$ is given by $K$ and $N_{G}(T)$ is the stabilizer of $AK$ (the calculation is done in the proof of 4.12). Thus $\operatorname{Stab}_{G}(e,AK)=N_{K}(T)$ follows and by 6.4 follows that $\widetilde{\mathcal{W}}M\cong N_{K}(T)$.
By 3.14 (A3) follows that $M$ is a subset of $K$ and centralizes $A$. Hence $M$ is a subgroup of $\operatorname{Fix}_{G}(p,F)$. Using 6.4 and the isomorphism Equation 10 it follows that
$\nicefrac{{\operatorname{Stab}_{G}(e,AK)}}{{M}}=\nicefrac{{N_{K}(T)}}{{M}}=\nicefrac{{\widetilde{\mathcal{W}}M}}{{M}}\cong\nicefrac{{\widetilde{\mathcal{W}}}}{{M^{\dagger}}}\cong\mathcal{W}.$
So far we have established the isomorphism $\operatorname{Stab}_{G}(e,AK)/M\cong\mathcal{W}$. Since the representation $\rho_{KM}\colon\mathcal{W}\to\operatorname{GL}(\mathfrak{a})$ (cp. Equation 7) is faithful, each element different from the identity of $\mathcal{W}$ acts non-trivially on $\mathfrak{a}$. This leads to the fact that $M=\operatorname{Fix}_{G}(p,F)$. ∎
Note that from these isomorphisms used in the previous proof it follows that any non-trivial coset of $\operatorname{Stab}_{G}(e,AK)/M$ acts non-trivially on the flat $AK$.
6.2 A Local to Global Approach
After we have discussed the local and global action of a Kac–Moody group on the corresponding symmetric space, we can try to connect them. In fact, one can do this using the Weyl groups and the fact that they are isomorphic, see 6.5. The following statement comes from a theorem given for non-affine real split Kac–Moody groups, see [FHHK20, Theorem 6.25]. Here we generalize it to Kac–Moody groups of arbitrary type, and since we use 6.5, our given assumption is that $G$ is a split real Kac–Moody group of arbitrary type $\mathbb{A}$, where $\mathbb{A}$ is non-spherical, indecomposable and symmetrizable. This enables us to use all the local action statements. Before we state the theorem, it should be noted that a Coxeter diagram can be obtained from a Dynkin diagram $\Gamma$ by forgetting the double edges. Therefore, each diagram automorphism of $\Gamma$ induces an automorphism of the corresponding Coxeter diagram.
In summary, the group $\operatorname{Aut}(\Gamma)$ embeds in $\operatorname{Aut}(\mathcal{W},S)$ and automorphisms of the embedded subgroup must preserve $S$ as a set.
Theorem 6.6
Let $G$ be a split real Kac–Moody group of arbitrary type $\mathbb{A}$, where $\mathbb{A}$ is non-spherical, indecomposable and symmetrizable. Further denote by $G^{\dagger}$ the centered subgroup of $G$ and with $X^{\dagger}=G^{\dagger}/K^{\dagger}$ the corresponding Kac–Moody symmetric space and let $(p,F)$ be a pointed maximal flat in $X^{\dagger}$. Then the following diagram commutes:
[…]
Here $\kappa$ and $\iota$ denotes canonical inclusion. Further, any local automorphism extends to a global automorphism if and only if $\operatorname{Aut}(\Gamma)=\operatorname{Aut}(\mathcal{W},S)$.
Proof:
The map $\iota$ is the natural inclusion since the Weyl group of the symmetric space takes values in the group of local transformations. The isomorphism of the right arrow follows directly from 6.3.
Now we are going to prove the isomorphism of the left arrow and there we use the same strategy as in the proof of Theorem 6.25 from [FHHK20]. Therefore, recall that
$W(\operatorname{Aut}(X^{\dagger})\curvearrowright X^{\dagger})\coloneqq\nicefrac{{\operatorname{Stab}_{\operatorname{Aut}(X^{\dagger})}(p,F)}}{{\operatorname{Fix}_{\operatorname{Aut}(X^{\dagger})}(p,F)}}.$
Since this definition of the Weyl group corresponding to the symmetric space is independent of the pointed flat (see [FHHK20, Proposition 2.36], which is proved for abstract symmetric spaces due to Loos), one can consider the standard pointed flat $(e,A^{\dagger}K^{\dagger})$.
Using the isomorphism given in 5.5, one can use the decomposition to check how the single parts acts on the standard flat. For this purpose, recall the explanation of the different automorphism types given in Section 5.1 or see [CM05, Section 8.2]. Furthermore, 5.5 yields the isomorphism
$\operatorname{Aut}(X^{\dagger})\cong\operatorname{Aut}_{\text{eff}}(G^{\dagger})=\left(\nicefrac{{G^{\dagger}}}{{C_{K^{\dagger}}(G^{\dagger})}}\rtimes(D\times\langle\Theta\rangle)\right)\rtimes\operatorname{Aut}(\Gamma).$
Now we use this decomposition of $\operatorname{Aut}(X^{\dagger})$ to determine the stabilizer and the fixator by considering each factor separately.
Since the diagram automorphism $\operatorname{Aut}(\Gamma)$ stabilizes the standard flat $(e,A^{\dagger}K^{\dagger})$ it remains in the decomposition and based on the fact that it is a subgroup of $\operatorname{Aut}(\mathcal{W},S)$, one has the inclusion $\kappa$ in this component. The diagonal automorphism fixes the standard flat pointwise and hence has no effect, i.e. the factor vanishes.
The Chevalley involution $\Theta$ acts on the standard flat by inversion and so it corresponds to $\mathbb{Z}/2\mathbb{Z}$. Note that the semidirect product with $\Theta$ vanishes in the decomposition of the stabilizer in $\operatorname{Aut}_{\text{eff}}(G^{\dagger})$ as one considers the action on $A^{\dagger}K^{\dagger}$ and by the proof of [FHHK20, Proposition 6.5] it is known that the Chevalley involution commutes with $\operatorname{Aut}(\Gamma)$ and it can be concluded that $\Theta$ also commutes with the inner automorphisms. In fact, since one considers the stabilizers of the standard flat $A^{\dagger}K^{\dagger}$, one know that $\Theta$ commutes with the diagram automorphism and normalizes the left multiplication. If we denote the left multiplication with $g\in\operatorname{Stab}_{G^{\dagger}}(e,A^{\dagger}K^{\dagger})$ by $l(g)$, then the following is true
$\Theta\circ l(g)=l(\Theta(g))\circ\Theta.$
After this, it is left to analyze the first factor in the decomposition of $\operatorname{Aut}(X^{\dagger})$. Note that the inner automorphism of $\operatorname{Aut}_{\text{eff}}(G^{\dagger})$ corresponds to $G^{\dagger}/C_{K^{\dagger}}(G^{\dagger})$, cp. 5.5. Further, by 6.5 one can consider the stabilizer and fixator of $A^{\dagger}K^{\dagger}$ as subgroups of $G^{\dagger}$:
$\operatorname{Stab}_{G^{\dagger}}(e,A^{\dagger}K^{\dagger})\cong N_{K^{\dagger}}(T^{\dagger})\quad\text{and}\quad\operatorname{Fix}_{G^{\dagger}}(e,A^{\dagger}K^{\dagger})\cong M^{\dagger}.$
Using 6.5 and the isomorphism theorem, the following can be calculated
/StabG(e,AK)/CK(G)FixG(e,AK)/CK(G) $\cong\nicefrac{{\operatorname{Stab}_{G^{\dagger}}(e,A^{\dagger}K^{\dagger})}}{{\operatorname{Fix}_{G^{\dagger}}(e,A^{\dagger}K^{\dagger})}}$
$\cong\mathcal{W}.$
So we have shown the isomorphisms of the vertical arrows, the commutativity of the diagram follows from the fact that the horizontal arrows in fact are inclusions. ∎
7 Comparing Symmetric Spaces
In the previous chapters we established a symmetric space associated to Kac–Moody groups over $\mathbb{K}\in\{\mathbb{R},\mathbb{C}\}$ and with respect to arbitrary generalized Cartan matrices. Now we can give an summary and compare it with the symmetric space developed in [FHHK20], which was set up for non-affine Kac–Moody groups over $\mathbb{R}$.
In [FHHK20] the authors consider only real centered Kac–Moody groups with respect to the root datum $\mathcal{D}_{sc}^{\mathbb{A}}$ (see [Mar18, Example 7.11]). This choice implies that their coroots are linearly independent, but this does not mean that the roots are linearly independent. For this reason, the authors in [FHHK20] work with the derived Kac–Moody algebra (see [Mar18, Definition 7.13]), and use the Cartan subalgebra of this Kac–Moody algebra to construct the standard flat of the associated symmetric space.
If we now consider Kac–Moody groups with repsect to arbitrary generalized Cartan matrices, one problem we encounter is that Cartan matrices do not have to be invertible. In the case of a non-invertible generalized Cartan matrix, there is a non-trivial kernel of bilinear form the Cartan subalgebra of the derived Kac–Moody algebra, which is equal to the center ([Kac90, Lemma 2.1 a)]).
To work around this problem for non-affine Cartan matrices, the authors of [FHHK20] use the semisimple adjoint quotient of a Kac–Moody group: this is the quotient of the group by the center of the Kac–Moody algebra that is lifted onto the group, cf. [FHHK20, Definition 3.8]. This quotient allows them to prove, for example, the useful statements [FHHK20, Proposition 6.4] and [FHHK20, Theorem 6.12 (ii)].
If we also consider Kac–Moody symmetric spaces over $\mathbb{C}$, we need a topological Iwasawa decomposition to define flats. At the time [FHHK20] was created, no complex version of the topological Iwasawa decomposition was known.
Here, we are talking about Kac–Moody groups over $\mathbb{R}$ and $\mathbb{C}$ with respect to the root datum $\mathcal{D}_{\mathrm{Kac}}^{\mathbb{A}}$, where $\mathbb{A}$ is an arbitrary generalized Cartan matrix. In this case the linear independence of the roots is given by enlarging the Cartan subalgebra, which leads to an enlarged torus at the group level.
This larger torus turns out to be a problem, because for the analysis of the global and local action of the Kac–Moody group on the corresponding symmetric space it is important that the two automorphism groups are related (see 5.5). The reason for this problem is Caprace’s theorem: It is the basis for understanding the automorphisms of a particular type of Kac–Moody group. But it only describes the automorphisms of a Kac–Moody group generated by its root groups, see [Cap09, Theorem 4.2]. On the extended torus one essentially asks for the automorphism group of $(\mathbb{R},+)$. This group is too big to classify, so it is not possible to say anything about the automorphisms on the torus outside of $G^{\dagger}$.
Therefore, in order to study the local and global action, this paper is restricted to the subgroup $G^{\dagger}$ of a Kac–Moody group $G$, which is generated only by its root groups. Caprace’s theorem applies to this centered subgroup $G^{\dagger}$ and leads to 5.5.
Regarding the case of generalizing the topological Iwasawa decomposition to complex split Kac–Moody groups, there is now a way to do this thanks to [HK23, Corollary B.8]. The quoted statement allows us to use the same methods as in [FHHK20] to prove the topological Iwasawa decomposition in the complex case, see 3.20.
So far, we have achieved a generalization of the Kac–Moody symmetric space described in [FHHK20].
8 Almost Split Kac–Moody Groups
After developing a theory of symmetric Kac–Moody spaces over $\mathbb{C}$ of arbitrary type $\mathbb{A}$, where $\mathbb{A}$ is an indecomposable, symmetrizable generalized Cartan matrix, it is of interest to ask about symmetric spaces of real forms of a Kac–Moody group. Thanks to the work of B. Rémy, [Rém02], there is a theory of real forms in the context of Kac–Moody groups. So in this section we first give a brief overview of the theory and the definitions of the necessary notions and then describe how to develop a theory of almost split Kac–Moody symmetric spaces. This section is mainly based on [Rém02].
8.1 Real Forms of Kac–Moody Groups
Following [Rém02, Chapter 11], we define several types of a real form.
Definition 8.1 (cf. Definition 11.1.2 from [Rém02])
[…]
  • A $\mathbb{R}$-functor $F$ is called a $\mathbb{R}$-form of a $\mathbb{C}$-functor $H$, if there is a functorial isomorphism $F\mid_{\mathbb{C}}\cong H$.
Clearly is the Tits functor $\mathcal{G}$ an example of an $\mathbb{R}$-form of itself.
Let $F$ be a $\mathbb{R}$-form of a $\mathbb{C}$-functor $H$. Now, we want to define an action of the Galois group $\Gamma\coloneqq\operatorname{Gal}(\mathbb{C}/\mathbb{R})$ on $H(\mathbb{C})$ using the real form $F$. Therefore denote by $R_{\mathbb{R}}$ a $\mathbb{R}$-algebra and by $R_{\mathbb{C}}=R_{\mathbb{R}}\otimes_{\mathbb{R}}\mathbb{C}$ the corresponding $\mathbb{C}$-algebra given by scalar extension. Then for all $\sigma\in\Gamma$ the map
$1_{R_{\mathbb{R}}}\otimes\sigma\colon R_{\mathbb{C}}\to R_{\mathbb{C}}$
is an isomorphism of $\mathbb{R}$-algebras. By the fact, that the the real form $F$ equals the $\mathbb{C}$-functor $H$ over $\mathbb{C}$, the isomorphism $1_{R_{\mathbb{R}}}\otimes\sigma$ induces an automorphism $f_{\sigma}$ on $H(R_{\mathbb{C}})$. This leads to the following commutative diagram.
[…]
To simplify the notation, we write $\sigma$ instead of $f_{\sigma}$. Note, that this action is a natural transformation, for more details see [Rém02, 11.1.2].
Definition 8.2
Let $G$ be a Kac–Moody group over $\mathbb{C}$ of type $\mathbb{A}$, where $\mathbb{A}$ is a indecomposable, symmetrizable generalized Cartan matrix. Further denote by $\mathcal{G}$ the Tits functor of type $\mathcal{D}_{\mathrm{Kac}}^{\mathbb{A}}$, i.e. $\mathcal{G}(\mathbb{C})=G$. Then call an $\mathbb{R}$-form $\mathcal{F}$ of the Tits functor $\mathcal{G}$ a functorial $\mathbb{R}$-form of the Kac–Moody group $G$.
Recall, that a Kac–Moody algebra $\mathfrak{g}(\mathbb{A})$ has an associated universal enveloping algebra $\mathcal{U}_{\mathfrak{g}(\mathbb{A})}$, see [Mar18, Chapter 3.1] or [Rém02, 7.3]. Following [Rém02, 7.4.5] we denote by $\mathcal{U}$ a corresponding $\mathbb{Z}$-form of $\mathcal{U}_{\mathfrak{g}(\mathbb{A})}$, i.e.
$\mathcal{U}\otimes_{\mathbb{Z}}\mathbb{C}\cong\mathcal{U}_{\mathfrak{g}(\mathbb{A})}.$
Denote by $\mathcal{U}_{\mathbb{R}}$ the scalar extension of $\mathcal{U}$ with respect to $\mathbb{R}$, i.e. $\mathcal{U}\otimes_{\mathbb{Z}}\mathbb{R}$. Note, that the universal enveloping algebra can be seen as a functor, which is very useful to define the notion of a prealgebraic $\mathbb{R}$-form:
The pair $(\mathcal{F},\mathcal{U}_{\mathbb{R}})$, where $\mathcal{F}$ is a functorial $\mathbb{R}$-form of a Kac–Moody group $G$ and $\mathcal{U}_{\mathbb{R}}$ is an $\mathbb{R}$-form of the universal enveloping algebra $\mathcal{U}_{\mathfrak{g}(\mathbb{A})}$, is called prealgebraic $\mathbb{R}$-form of $G$under the following conditions:
[…]
  • For each subfield $\mathbb{K}$ of $\mathbb{C}$ is $\mathcal{F}(\mathbb{K})$ a subgroup of $\mathcal{G}(\mathbb{C})=G$.
For more details about prealgebraic $\mathbb{R}$-froms see [Rém02, 11.1.3] and for more details about the adjoint representation see [Rém02, 9.5].
From [Rém02, Proposition 11.2.2] it follows that the Galois action on a Kac–Moody group $G$ stabilizes the set of tori in $G$ and that for every Galois action $\sigma$ there is an element $g\in G$, so that one can rectify the action in the sense that the standard torus is fixed. In detail, this means that the action
$c_{g^{-1}}\circ\sigma$
is well-defined and fixes the standard torus $T\leq G$. In the following denote by $\overline{\sigma}$ the rectification of $\sigma$.
Since a root $\alpha\in\Delta^{\mathrm{re}}$ is defined with respect to the Cartan subalgebra $\mathfrak{h}$, see Section 3.1, one can speak of a root $\alpha$ relative to $\mathfrak{h}$. In this context, we can also call a root group $U_{\alpha}$ relative to $T$ if the root $\alpha$ is relative to the Cartan subalgebra $\mathfrak{h}$ underlying $T$.
Convention 8.3 (cf. Hypothesis 11.2.3 from [Rém02].)
Denote by (SGR) the condition, that a rectified Galois action of a Kac–Moody group maps a root group relative to $T$ to a root group of the same type.
Condition (SGR) induces an action of the rectified Galois action on the set of real roots $\Delta^{\mathrm{re}}$. To make use of this action, we give the following definition, cf. [Rém02, 11.2.5].
Definition 8.4
A prealgebraic $\mathbb{R}$-form $(\mathcal{F},\mathcal{U}_{\mathbb{R}})$ of a Kac–Moody group $G$ over $\mathbb{C}$ of type $\mathbb{A}$, where $\mathbb{A}$ is an indecomposable, symmetrizable generalized Cartan matrix, is called an algebraic $\mathbb{R}$-form of $G$ if
  • the condition (SGR) is satisfied,
  • for every rectified Galois action $\overline{\sigma}$ (i.e. $\overline{\sigma}(T)=T$) the following is true
    • the action $\overline{\sigma}$ respects the decomposition into $Q$-homogeneous components of $\mathcal{U}_{\mathbb{R}}$ and the induced permutation on $Q$ satisfies the following homogeneity condition:
      For all $\alpha\in\Delta^{\mathrm{re}}$ and for all $n\in\mathbb{N}$ holds: $\overline{\sigma}(n\alpha)=n\overline{\sigma}(\alpha)$.
    • the action $\overline{\sigma}$ stabilizes $\Lambda$ and $\Lambda^{\vee}$.
Furthermore, if the Galois action stabilizes the conjugacy class of the positive Borel subgroups of $G$, call the algebraic $\mathbb{R}$-form an almost split Kac–Moody $\mathbb{R}$-group.
Note that by [Rém02, Lemma 11.3.2] one can rectify a Galois automorphism $\sigma$ in such a way that the sign of the Borel subgroup is preserved. Moreover, this rectification also stabilizes the standard torus $T$ and is well-defined up to $T$. Thus, this rectification is just a special case of the rectified Galois automorphism we defined above. In particular, this action induces an action on the Weyl group $\mathcal{W}$ and stabilizes the set of generators $S$, see [Rém02, Proposition 11.3.2]. Since we are only interested in almost split Kac–Moody $\mathbb{R}$ groups and for simplicity, we denote this rectification by $\overline{\sigma}$. For more details about almost split Kac–Moody $\mathbb{R}$-groups see [Rém02, 11.3.1].
Before we discuss Galois descent, we give useful properties of the rectified Galois action. Let $G$ be a Kac–Moody group over $\mathbb{C}$ of type $\mathbb{A}$, where $\mathbb{A}$ is indecomposable and symmetrizable. Further let $(\mathcal{F},\mathcal{U}_{\mathbb{R}})$ be a algebraic $\mathbb{R}$-form of $G$. Then any rectified Galois automorphism $\overline{\sigma}$ has the following properties:
  1. 1
    For any real root $\alpha\in\Delta^{\mathrm{re}}$ and any parameter $t\in\mathbb{C}$, there is the equation
    $\overline{\sigma}(x_{\alpha}(t))=x_{\overline{\sigma}(a)}(k_{\alpha}\sigma(t)),$
    where $k_{\alpha}\in\mathbb{C}^{\times}$ is an element defined by the Galois action on a generator of the Kac–Moody algebra.
  2. 2
    The reflection $r_{\alpha}\in\mathcal{W}$ is mapped to $r_{\overline{\sigma}(\alpha)}$.
  3. 3
    For any real root $\alpha\in\Delta^{\mathrm{re}}$ the element $h_{\alpha}\in\Lambda^{\vee}$ is mapped to $h_{\overline{\sigma}(\alpha)}$.
For details we refer to [Rém02, Proposition 11.2.5].
8.2 The Galois Descent
In linear algebra it is well known that a real form of a complex vector space can be identified with the fixed point set of an action given via Galois automorphism. The same idea uses Rémy to develop further the theory of almost split Kac–Moody $\mathbb{R}$-groups.
In [Rém02, Chapter 12], he uses the fact that one can associated to a Kac–Moody group a twin building on which he can also define a Galois action which is compatible with the Galois automorphism action on the group. In detail, he uses a geometric realization to study the Galois action and for example to check if there exists at all fixed points, for the realization see [Rém02, Chapter 5] and for the action on the realization of the twin building see [Rém02, 12.1.2].
The theory Rémy develops in Chapter 12 is very technical and uses a lot of new notions, which we do not need in detail, hence, we only give a short overview. First let $G$ be a Kac–Moody group over $\mathbb{C}$ of type $\mathbb{A}$, where $\mathbb{A}$ is an indecomposable, symmetrizable generalized Cartan matrix. Assume that $G$ has an algebraic $\mathbb{R}$-form $(\mathcal{F},\mathcal{U}_{\mathbb{R}})$ which satisfies (SGR), i.e. $G$ is an almost split Kac–Moody $\mathbb{R}$-group. Then call
$G(\mathbb{R})\coloneqq\mathcal{F}(\mathbb{R})$
the rational points of $G$. Furthermore, $G$ satisfies the descent condition (8.2) if
$G(\mathbb{R})=G^{\Gamma},$
where $\Gamma$ denotes the group of the induced Galois automorphism on $G$, see Section 8.1. Call a subgroup $H$ of $\mathcal{F}(\mathbb{C})=G$ an $\mathbb{R}$-subgroup, if it is stable under the Galois action. The rational points of an $\mathbb{R}$-subgroup $H\leq G$, denoted by $H(\mathbb{R})$ are given by the Galois fixed points of $H$.
To make the following terms well-defined, we need to define a triple $(A,C,-C)$, where $A$ is an apartment and $C,-C$ are opposite chambers. This is similar to defining a standard torus and defines the geometric realization. In our context, it is only necessary that we obtain the following when applying the theory of [Rém02, Chapter 12]:
  • The group of rational points possess a root system, which is not necessarily reduced. A real $\mathbb{R}$-root is a restricted root in the context of the geometric realization. To be precise, a root (there it is called half-apartment) in the geometric realization is given by a hyperplane on which a reflection is defined. In the process of doing the Galois descent, one restricts the hyperplane to the fixed point set. For more details see [Rém02, Chapter 5]. There the geometric realization and the relative terminology is defined.
    Denote by $\musNaturalre\Delta^{\musNatural}_{\mathrm{re}}$ the set of all real $\mathbb{R}$-roots with respect to $(A,C,-C)$ (cf. [Rém02, 5.5.1]). Denote by $\musNatural\Pi^{\musNatural}$ the set of simple roots of $\musNaturalre\Delta^{\musNatural}_{\mathrm{re}}$.
[…]
  • By [Rém02, Theorem 12.6.3] it follows, that the group of rational points $G(\mathbb{R})$ of an almost split Kac–Moody $\mathbb{R}$-group $G$ possess an RGD system.
8.3 Relative Kac–Peterson Topology
As the Kac–Peterson topology is defined for a Kac–Moody group $G$, see Section 3.2 for the definition, we can define a relative Kac–Peterson topology on the rational points $G(\mathbb{R})\leq G$ of an almost split Kac–Moody $\mathbb{R}$-group. The question is whether the relative Kac–Peterson topology coincides with the subspace topology induced by the surrounding Kac–Moody group. This section is based on [Voc23, 8.2.3].
[…]
where $\musNatural\alpha_{i}\in\Pi^{\musNatural}$.
Even the torus $Z(\mathbb{R})\leq G(\mathbb{R})$, defined in [Rém02, Chapter 12.5], is a subgroup of the torus of the surrounding split Kac–Moody group and splits over $\mathbb{R}$, see [Rém02, 12.5.1] and [Rém02, 12.5.3]. This means that, that $Z(\mathbb{R})$ also carries a Lie group topology.
Now we have all the ingredients to define the relative Kac–Peterson topology in the same way as the Kac–Peterson topology, comparing to Section 3.2.
Definition 8.5
Let $G$ be an almost split Kac–Moody $\mathbb{R}$-group satisfying (8.2) and denote by $G(\mathbb{R})$ the rational points of $G$. Then the relative Kac–Peterson topology $\tau_{\text{KP}}^{\operatorname{as}}$ on $G(\mathbb{R})$ is the finest group topology on $G(\mathbb{R})$ such that the embeddings of the relative rank one groups and the embedding of $Z(\mathbb{R})$ are continuous with respect to their induced Lie topologies.
In the proof of [HK23, Porposition B.3] bounded subgroups are used to define a topology on a split Kac–Moody group $G$ which euqals the Kac–Peterson topology. This can also be done for rational points of an almost split Kac–Moody $\mathbb{R}$-group $G(\mathbb{R})$. To define the notion of a restricted subgroup in $G(\mathbb{R})$, we recall what a parabolic subgroup is in the context of a split Kac–Moody group $G$ over $\mathbb{C}$, see [Mar18, p.150]:
Denote by $\mathcal{W}_{J}$ the subgroup of the Weyl group $\mathcal{W}$ given by $\mathcal{W}_{J}=\langle r_{i}\mid i\in J\rangle$ for a subset $J\subseteq I$, where $I=\{1,\ldots,n\}$ , see [AB08, Definition 2.12]. Then a parabolic subgroup $P_{J}$ is defined by
$P_{J}=B^{+}\mathcal{W}_{J}B^{+}.$
If $\mathcal{W}_{J}$ is finite, call $P_{J}$ spherical. In the language of buildings, a parabolic subgroup is the stabilizer in $G$ of the standard residue $J$ in the positive half of the twin building $\Delta_{\pm}$.
Definition 8.6
Let $G$ be a split Kac–Moody group over $\mathbb{K}\in\{\mathbb{R},\mathbb{C}\}$ of type $\mathbb{A}$, where $\mathbb{A}$ is an indecomposable, symmetrizable generalized Cartan matrix. Call a subgroup $M\leq G$ bounded if it is contained in the intersection of two spherical parabolic subgroups with opposite signs.
Since bounded subgroups are subgroups contained in spherical parabolic subgroups, they carry a Lie group structure and thus a Lie group topology.
Definition 8.7
Let $G$ be a split Kac–Moody group over $\mathbb{K}\in\{\mathbb{R},\mathbb{C}\}$ of type $\mathbb{A}$, where $\mathbb{A}$ is an indecomposable, symmetrizable generalized Cartan matrix. Denote by $\operatorname{MBS}(\mathbb{K})$ the set of maximal bounded subgroups. The maximal bounded subgroup topology $\tau_{\text{MB}}$ is the finest group topology, so that the embeddings of the maximal bounded subgroups in $G$ are continuous.
Note, that [HK23, Proposition B.3] is formulated for centered split Kac–Moody groups, thus we now prove a version for arbitrary split Kac–Moody groups.
Proposition 8.8
Let $G$ be a split Kac–Moody group over $\mathbb{K}\in\{\mathbb{R},\mathbb{C}\}$ of type $\mathbb{A}$, where $\mathbb{A}$ is an indecomposable, symmetrizable generalized Cartan matrix. Equip the standard torus with the Lie topology induced by the exponential map, see 3.7. Then the Kac–Peterson topology $\tau_{\text{KP}}$ equals the maximal bounded subgroup topology $\tau_{\text{MB}}$.
Proof:
We follow the strategy of the proof given for [HK23, Proposition B.3]. Since the cited proposition is given for centered Kac–Moody groups and in general a Kac–Moody group $G$ can be decomposed in $G=TG^{\dagger}$, see 3.10, where $G^{\dagger}$ is a centered Kac–Moody group, see 3.13, we has to take a closer look at the tours.
Note, that by definition the torus is a bounded subgroup,
$T=B_{+}\cap B_{-},$
and carries a Lie group topology, see 3.7. Thus the torus embeds as a closed subgroup into a maximal bounded subgroup and by the closed-subgroup theorem ([HN12, Theorem 9.3.7]) follows that the subspace topology of the embedded torus equals its Lie group topology.
As a next step, we apply [Mar15, Lemma 4.3], which states that the subspace topology induced by $\tau_{\text{KP}}$ on the maximal bounded subgroups is equal to the Lie group topology. Although the statement is given for centered split Kac–Moody groups, it also applies to all split Kac–Moody groups, since the statement is based on building theory arguments.
Recall, that the Kac–Peterson topology induces the Lie group topology on the rank one subgroups $G_{\alpha}\cong\operatorname{SL}_{2}(\mathbb{K})$, $\alpha\in\Delta^{\mathrm{re}}$, see [HK23, Definition and Remark 2.1]. Furthermore, the rank one subgroups are bounded subgroups and a rank one subgroup embeds as a closed subgroups into a maximal bounded subgroup. By the closed-subgroup theorem ([HN12, Theorem 9.3.7]) follows that the subspace topology of the embedded rank one subgroups in a maximal bounded subgroup equals its Lie group topology.
Overall it follows that
$\tau_{\text{MB}}\subseteq\tau_{\text{KP}}.$
To see the other direction, consider the maximal bounded subgroups in $(G,\tau_{\text{KP}})$. Here, each maximal bounded subgroup has the Lie group topology induced by the subspace topology (cf. [Mar15, Lemma 4.3]). The embedded relative rank one subgroups and the embedded torus also have the induced subspace Lie topology and one concludes that the embedding of the maximal bounded subgroups in $G$ is continuous with respect to the topology $\tau_{\text{KP}}$. Therefore, the topology $\tau_{\text{KP}}$ is coarser than $\tau_{\text{MB}}$. ∎
A parabolic $\mathbb{R}$-subgroup of $G(\mathbb{R})$ is also a parabolic subgroup in the surrounding split Kac–Moody group $G$, see [Rém02, Definition 12.1.3] and [Rém02, Remark (1) 12.1.3]). Thus a maximal bounded $\mathbb{R}$-subgroup embeds in a maximal bounded subgroup of $G$. By [Rém02, 6.2] and [Cap09, Section 3.1] follows that maximal bounded subgroups are (algebraic) Lie groups. To keep the notation simple, we denote the set of maximal bounded subgroups in $G(\mathbb{R})$ by $\operatorname{MBS}(\mathbb{R})$ and in $G$ by $\operatorname{MBS}(\mathbb{C})$.
[…]
From the closed-subgroup Theorem (cf. [HN12, Theorem 9.3.7]) and the fact that a relative rank one subgroup is a closed subgroup in a maximal bounded subgroup of $\operatorname{MBS}(\mathbb{C})$ because the Galois action on $G$ is continuous, it follows that a relative rank one subgroup is an embedded Lie group. Moreover, the Lie group topology coincides with the subspace topology from $\operatorname{MBS}(\mathbb{C})$. Note that the subspace topology induced by $\tau_{\text{KP}}$ on a maximal bounded group of $\operatorname{MBS}(\mathbb{C})$ is equal to the Lie group topology, see the proof of 8.8.
Now, we consider the topology that can be defined by the embedding of the maximal bounded subgroups. Recall the definition of $\tau_{\text{MBS}}$, see 8.7.
Definition 8.9
Let $G$ be an almost split Kac–Moody $\mathbb{R}$-group satisfying (8.2) and denote by $G(\mathbb{R})$ the rational points of $G$. Define the topology $\tau_{\text{MB}}^{\operatorname{as}}$ on $G(\mathbb{R})$ as the finest group topology so that the embeddings of the maximal bounded $\mathbb{R}$-subgroups are continuous.
Lemma 8.10
Let $G$ be an almost split Kac–Moody $\mathbb{R}$-group satisfying (8.2) and denote by $G(\mathbb{R})$ the rational points of $G$. Then
$\tau_{\text{MB}}^{\operatorname{as}}=\tau_{\text{MB}}\mid_{G(\mathbb{R})},$ (11)
where $\tau_{\text{MB}}\mid_{G(\mathbb{R})}$ means the subspace topology induced by $\tau_{\text{MB}}$.
Proof:
Both topologies are defined as the finest topologies, moreover they coincide on the $\operatorname{MBS}(\mathbb{R})$:
Every parabolic $\mathbb{R}$-subgroup of $G(\mathbb{R})$ is a parabolic subgroup of $G$. From the previous discussion, we know that the maximal bounded subgroups are Lie groups. This leads to the fact that $\operatorname{MBS}(\mathbb{R})$ are also Lie groups, and in particular they are equal to $\operatorname{MBS}(\mathbb{C})\cap G(\mathbb{R})$, equipped with the trace topology. ∎
In the last step, we want to compare the topologies. To do this, we need a different point of view for the Kac–Peterson topology:
The universal topology $\tau_{\text{Prod}}$, defined in [HKM13, Definition 7.19] using colomits, is equal to the Kac–Peterson topology (cp. [HKM13, Proposition 7.21]) and by transport of structure, it is a $k_{\omega}$ topology, for details see Section 3.6 or [FT77]. The authors of the article [GGH10, Corollary 6.7] states that $G(\mathbb{R})$ is $k_{\omega}$ with respect to the induced subspace topology coming from the surrounding split Kac–Moody group.
Putting all things from this section together like a puzzle, we observe that by 8.8 $\tau_{\text{KP}}$ is equal to $\tau_{\text{MB}}$ for split Kac–Moody groups. Moreover, by [HKM13, Proposition 7.21], one knows that $\tau_{\text{Prod}}=\tau_{\text{KP}}$. In particular
$\tau_{\text{MB}}=\tau_{\text{KP}}=\tau_{\text{Prod}}.$ (12)
This can be now used to compare the different structures on $G(\mathbb{R})$.
Theorem 8.11
Let $G$ be an almost split Kac–Moody $\mathbb{R}$-group satisfying (8.2) and denote by $G(\mathbb{R})$ the rational points of $G$. Then the following holds for $G(\mathbb{R})$:
  1. I.
    $\tau_{\text{MB}}^{\operatorname{as}}=\tau_{\text{MB}}\mid_{G(\mathbb{R})}=\tau_{\text{Prod}}\mid_{G(\mathbb{R})}=\tau_{\text{KP}}\mid_{G(\mathbb{R})}$,
  2. II.
    $G(\mathbb{R})$ is a $k_{\omega}$-space with respect to the topology $\tau_{\text{MB}}^{\operatorname{as}}$, and
  3. III.
    $\tau_{\text{KP}}^{\operatorname{as}}=\tau_{\text{MB}}^{\operatorname{as}}$.
Proof:
  1. I.:
    This statement follows from the discussion above with Equation 11.
  2. II.:
    This statement follows from the previous discussion.
  3. III.:
    Denote by $(\operatorname{MBS}(\mathbb{R})_{i})_{i\in I}$ the family of maximal bounded $\mathbb{R}$-subgroups in $G(\mathbb{R})$ and consider the topological spaces $(G(\mathbb{R}),\tau_{\text{KP}}^{\operatorname{as}})$ and $(G(\mathbb{R}),\tau_{\text{MB}}^{\operatorname{as}})$. A relative rank one subgroup embeds as a Lie subgroup into one of the $\operatorname{MBS}(\mathbb{R})_{i}$. Moreover the continuous embeddings of the relative rank one subgroups into $G(\mathbb{R})$ provide fewer conditions than the embeddings of the maximal bounded subgroups. It follows that the map from $(G(\mathbb{R}),\tau_{\text{KP}}^{\operatorname{as}})$ to $(G(\mathbb{R}),\tau_{\text{MB}}^{\operatorname{as}})$ is continuous, i.e. $\tau_{\text{MB}}^{\operatorname{as}}\subseteq\tau_{\text{KP}}^{\operatorname{as}}.$
    For the other direction consider the maximal bounded subgroups in the topological group $(G(\mathbb{R}),\tau_{\text{KP}}^{\operatorname{as}})$. Every $\operatorname{MBS}(\mathbb{R})_{i}$ carries the Lie group topology and the embedded relative rank one subgroups carry the induced Lie subspace topology. Thus one deduces that the embedding of $\operatorname{MBS}(\mathbb{R})_{i}$ into $G(\mathbb{R})$ with respect to the topology $\tau_{\text{KP}}^{\operatorname{as}}$ is continuous. To be more precise, the topology $\tau_{\text{KP}}^{\operatorname{as}}$ comes from the embedding of relative rank one subgroups equipped with the Lie topology and by the closed-subgroup theorem (cf. [HN12, Theorem 9.3.7]) the maximal bounded subgroups carries the same Lie group topology. This means that the topology $\tau_{\text{KP}}^{\operatorname{as}}$ is coarser than $\tau_{\text{MB}}^{\operatorname{as}}$.
9 Symmetric Spaces of Almost Split Kac–Moody Groups
In this section, we going to define a symmetric space for almost split Kac–Moody $\mathbb{R}$-groups.
Therefore, let $G$ be a Kac–Moody group over $\mathbb{C}$ of type $\mathbb{A}$, where $\mathbb{A}$ is an indecomposable, symmetrizable generalized Cartan matrix of size $n$ with rank $l$. Denote by $\Gamma$ the group of Galois automorphism induced by $\operatorname{Gal}(\mathbb{C}/\mathbb{R})$ on $G$ and assume that $G$ possess an algebraic $\mathbb{R}$-form $(\mathcal{F},\mathcal{U}_{\mathbb{R}})$ satisfying the (8.2) condition, i.e.
$G(\mathbb{R})=G^{\Gamma},$
where $G(\mathbb{R})$ denotes the rational points of the almost split Kac–Moody $\mathbb{R}$-group $G$. Denote by $\overline{\sigma}$ the rectified Galois automorphism which stabilizes the standard torus $T$ of $G$, i.e. $\overline{\sigma}(T)=T$.
Lemma 9.1
Let $G$ be an almost split Kac–Moody $\mathbb{R}$-group. Any rectified Galois automorphism $\overline{\sigma}$ can be conjugated by a diagonal automorphism such that it commutes with $\Theta$.
Proof:
Similar to the proof of the algebraic Iwasawa decomposition, we use the decomposition
$G=TG^{\dagger}.$
This allows us to prove the statement separately for $T$ and $G^{\dagger}$. Note, that by Section 5.1 every automorphism of $\operatorname{Aut}(G^{\dagger})$ can be decomposed into an inner automorphism, a sign automorphism, a diagonal automorphism, a diagram automorphism and a field automorphism.
By Section 8.1 1 one knows how $\overline{\sigma}$ acts on an element of a root group $U_{\alpha}$, $\alpha\in\Delta^{\mathrm{re}}$,:
$\overline{\sigma}(x_{\alpha}(t))=x_{\overline{\sigma}(\alpha)}(k_{a}\overline{t}),$
where $x_{\alpha}(t)$, $t\in\mathbb{C}$, denotes an element $U_{\alpha}$, see Section 3.2. The operation to add $k_{\alpha}$ can be understood as the action of a diagonal automorphism. Hence $\overline{\sigma}$ can be rectified by conjugation with a diagonal automorphism $d$ so that $k_{\alpha}$ becomes $1$. In detail denote by
$\widehat{\sigma}\coloneqq d\circ\overline{\sigma}\circ d^{-1}$
the rectified Galois automorphism, where $d$ acts by $d(x_{\alpha}(t))=x_{\alpha}(d_{\alpha}t)$ such that $k_{\alpha}=\frac{\overline{d_{\alpha}}}{d_{\overline{\sigma}(a)}}$. This results in
$\widehat{\sigma}(x_{\alpha}(t))$ $=(d\circ\overline{\sigma}\circ d^{-1})(x_{\alpha}(t))$
$=(d\circ\overline{\sigma})(x_{\alpha}(d_{\alpha}^{-1}t))$
$=d(x_{\overline{\sigma}(\alpha)}(k_{\alpha}\overline{d_{\alpha}^{-1}}\overline{t}))$
$=x_{\overline{\sigma}(\alpha)}(d_{\overline{\sigma}(\alpha)}\overline{d_{\alpha}^{-1}}k_{\alpha}\overline{t})$
$=x_{\overline{\sigma}(\alpha)}(\overline{t}).$
By [CM05, p.384] one can always choose such a diagonal automorphism with respect to $d_{\alpha}$ and $d_{\overline{\sigma}(\alpha)}$. Moreover, $\widehat{\sigma}(T)=T$, since the diagonal automorphism acts as the identity on $T$ by definition (cf. [CM05, p.384]).
This results in the observation, that $\Theta$ and $\widehat{\sigma}$ commutes on $G^{\dagger}$:
$\left(\Theta\circ\widehat{\sigma}\right)(x_{\alpha}(t))$ $=\Theta(x_{\overline{\sigma}(\alpha)}(\overline{t}))$
$=x_{-\overline{\sigma}(\alpha)}(t),$
and
$\left(\widehat{\sigma}\circ\Theta\right)(x_{\alpha}(t))$ $=\widehat{\sigma}\left(x_{-\alpha}(\overline{t})\right)$
$=x_{-\overline{\sigma}(\alpha)}(t).$
[…]
$\overline{\sigma}\left(r^{v_{i}}\right)=\overline{r}^{v_{\overline{\sigma}(i)}}.$
In Section 3.5 we have defined the action of $\Theta$ on $T$ and hence we can compute
$\left(\Theta\circ\overline{\sigma}\right)\left(r^{v_{i}}\right)$ $=\Theta\left(\overline{r}^{v_{\overline{\sigma}(i)}}\right)=r^{-v_{\overline{\sigma}(i)}},$
and on the other hand
$\left(\overline{\sigma}\circ\Theta\right)\left(r^{v_{i}}\right)$ $=\overline{\sigma}\left(\overline{r}^{-v_{i}}\right)=r^{-v_{\overline{\sigma}(i)}}.$
Based on the previous lemma, we observe that
$\Theta\left(G^{\Gamma}\right)=\Theta(G)^{\Gamma},$
where we denote for simplicity reasons with $\Gamma$ the group of rectified and conjugated Galois automorphism $\widehat{\sigma}$.
It follows that $\Theta$ is a continuous involution on the topological group $G(\mathbb{R})$ and we define
$K(\mathbb{R})\coloneqq K^{\Gamma}=\left(G^{\Theta}\right)^{\Gamma}=\left(G^{\Gamma}\right)^{\Theta}.$
Similar to the approach for split Kac–Moody groups, we can define the twist map on $G(\mathbb{R})$ by
$\tau^{\operatorname{as}}\colon G(\mathbb{R})$ $\to G(\mathbb{R})$
$g$ $\mapsto g\Theta\left(g^{-1}\right),$
and also a reflection map by
$\begin{split}\mu^{\operatorname{as}}\colon G(\mathbb{R})/K(\mathbb{R})\times G(\mathbb{R})/K(\mathbb{R})&\to G(\mathbb{R})/K(\mathbb{R})\\ (gK(\mathbb{R}),hK(\mathbb{R}))&\mapsto\tau^{\operatorname{as}}(g)\Theta(h)K(\mathbb{R}).\end{split}$ (13)
[…]
$\tau^{\operatorname{as}}(G(\mathbb{R}))\cap K(\mathbb{R})\subseteq\tau(G)\cap K=\{e\}.$
This results in the following theorem, which is a generalization of the split case.
Theorem 9.2
Let $G$ be an almost split Kac–Moody $\mathbb{R}$-group satisfying (8.2) and denote by $G(\mathbb{R})$ the rational points of $G$. The quotient
$X(\mathbb{R})\coloneqq\nicefrac{{G(\mathbb{R})}}{{K(\mathbb{R})}}$
is a symmetric space in the sense of Loos with the reflection map $\mu^{\operatorname{as}}$. Moreover, there is a natural action of $G(\mathbb{R})$ on the symmetric space given by automorphisms
$G(\mathbb{R})$ $\to\operatorname{Sym}(X(\mathbb{R}))$
$g$ $\mapsto(hK(\mathbb{R})\mapsto ghK(\mathbb{R})).$
Since $\mathcal{F}(\mathbb{C})=G$ if $G$ is a complex (split) Kac–Moody group, it follows that the theorem is in this particular case just 4.4 for $\mathbb{K}=\mathbb{C}$. Furthermore, if we consider a real split Kac–Moody group in the theorem, the following theorem corresponds to 4.4 for $\mathbb{K}=\mathbb{R}$.
Proof:
Since the properties of an abstract symmetric space due to Loos are satisfied, we need to check that
$G(\mathbb{R})\to\operatorname{Sym}(X(\mathbb{R}))\,,\;g\mapsto(hK(\mathbb{R})\mapsto ghK(\mathbb{R}))$
is an action. Thus, let $g,h,d\in G(\mathbb{R})$ and do the following calculation
$\mu^{\operatorname{as}}(ghK(\mathbb{R}),gdK(\mathbb{R}))$ $=\tau^{\operatorname{as}}(gh)\Theta(gd)K(\mathbb{R})$
$=gh\Theta(gh)^{-1}\Theta(g)\Theta(d)K(\mathbb{R})$
$=gh\Theta(h)^{-1}\Theta(g)^{-1}\Theta(g)\Theta(d)K(\mathbb{R})$
$=g(\tau^{\operatorname{as}}(h)\Theta(d)K(\mathbb{R}))$
$=g\mu^{\operatorname{as}}(hK(\mathbb{R}),dK(\mathbb{R})).$
9.1 Flats in Almost Split Kac–Moody Symmetric Spaces
After establishing the notion of a symmetric space for almost split Kac–Moody groups, we can define flats. As we in the proof of 4.8, we define a symmetric Euclidean space in the fixed point set of the Cartan subalgebra and use the exponential map to obtain an isomorphism.
Let $G$ be an almost split Kac–Moody $\mathbb{R}$-group. Then, based on Section 8.1 3, we know that the Galois action on $G$ also acts on $\mathfrak{h}$. Further, a rectified Galois action stabilizes the standard torus $T$ of $G$ and since $G$ satisfies the condition (8.2), we can define
[…]
The next step is to analyze the image of $\musNatural\exp_{\operatorname{KM}}(\mathfrak{a}^{\musNatural})$. Therefore, let
$h=\sum_{i=1}^{2n-l}r_{i}v_{i}\in\mathfrak{h}=\Lambda^{\vee}\otimes_{\mathbb{Z}}\mathbb{C}$
be an element of the Cartan algebra, where $r_{i}\in\mathbb{C}$ are scalars and $v_{i}\in\Lambda^{\vee}$ are the $\mathbb{Z}$-basis of $\Lambda^{\vee}$ (cp. Section 3.2). Now we can compute
$\overline{\sigma}(\exp_{\operatorname{KM}}(h))$ $=\overline{\sigma}\left(\prod_{i=1}^{2n-l}\left(e^{r_{i}}\right)^{v_{i}}\right)=\prod_{i=1}^{2n-l}\sigma(e^{r_{i}})^{v_{\overline{\sigma}(i)}}=\prod_{i=1}^{2n-l}(e^{\overline{r_{i}}})^{v_{\overline{\sigma}(i)}}$
$\exp_{\operatorname{KM}}(\overline{\sigma}(h))$ $=\exp_{\operatorname{KM}}\left(\sum_{i=1}^{2n-l}\overline{r_{i}}v_{\overline{\sigma}(i)}\right)=\prod_{i=1}^{2n-l}(e^{\overline{r_{i}}})^{v_{\overline{\sigma}(i)}}.$
The image of the exponential map restricted to $\musNatural\mathfrak{a}^{\musNatural}$ is given by $A(\mathbb{R})\coloneqq A\cap G(\mathbb{R})$, since the exponential map is Galois-equivariant.
Note that $\mathfrak{a}^{\musNatural}\cong\mathbb{R}^{m}$, where $m<2n-l$, but the actual size of $m$ is not clear. It depends on how exactly the torus of $G(\mathbb{R})$ looks and splits.
Proposition 9.3
Let $X(\mathbb{R})=G(\mathbb{R})/K(\mathbb{R})$ be the symmetric space attached to the group of rational points of an almost split Kac–Moody $\mathbb{R}$-group $G$. Moreover, equip $\musNatural\mathfrak{a}^{\musNatural}$ with the Euclidean symmetric space structure. Then for any $g\in G(\mathbb{R})$ there is an isomorphism of symmetric spaces given by
[…]
The subset $gA(\mathbb{R})K(\mathbb{R})$ of $X(\mathbb{R})$ is a Euclidean flat of dimension $\musNatural\dim\mathfrak{a}^{\musNatural}$. Call for any $g\in G(\mathbb{R})$ the set $gA(\mathbb{R})K(\mathbb{R})$ a standard $\mathbb{R}$-flat.
Proof:
Since $\varphi^{\operatorname{as}}_{g}$ is obtained from the map of 4.8 by restricting the exponential map to a subset of $\mathfrak{a}$, the assertion that $\varphi^{\operatorname{as}}_{g}$ is an isomorphism of symmetric spaces follows with the same arguments as in the proof of 4.8.
Assume that $g=e$, then it remains to check that $A(\mathbb{R})K(\mathbb{R})$ is a closed subset, since the multiplication is continuous. To do this, we can use the Iwasawa decomposition on $G$ and the fact, that $AK$ is a closed subset in the symmetric space associated to $G$. By 8.11, the induced subspace topology on $G(\mathbb{R})$ is the same as the relative Kac–Peterson topology on $G(\mathbb{R})$. Hence, if $AK\cap G(\mathbb{R})=A(\mathbb{R})K(\mathbb{R})$, the desired statement follows.
Recall from 3.20 that an element $g\in AK$ can be written as $ak$ for a unique $a\in A$ and a unique $k\in K$. Assume that $ak\in AK\cap G(\mathbb{R})\subseteq G$ and let $\overline{\sigma}$ a rectified Galois automorphism. Then it follows that $\overline{\sigma}(A)\subseteq A$ and $\overline{\sigma}(K)\subseteq K$. From the condition (8.2) it further follows that
$\overline{\sigma}(ak)=ak=\overline{\sigma}(a)\overline{\sigma}(k).$
By the uniqueness of the decomposition, $AK\cap G(\mathbb{R})=A(\mathbb{R})K(\mathbb{R})$ follows. Hence the set $A(\mathbb{R})K(\mathbb{R})$, and so $gA(\mathbb{R})K(\mathbb{R})$ for all $g\in G(\mathbb{R})$, is closed in $X(\mathbb{R})$. ∎
As for symmetric spaces of split Kac–Moody groups, we can give the characterization of flats. This is possible, since almost split Kac–Moody $\mathbb{R}$-groups posses an RGD system and the root groups has a good algebraic group structure.
In the proof of the characterization, one uses that the group model of the symmetric spaces is isomorphic to the coset model. But, the isomorphism between the coset model and the group model established in Section 4.3 is independent of the Kac–Moody group. It is essentially based on the isomorphism specified in 2.4 for abstract symmetric spaces according to Loos and on the existence of the twist map $\tau$. Hence, it follows that
$(X(\mathbb{R}),\mu^{\operatorname{as}})\cong(\tau^{\operatorname{as}}(G(\mathbb{R})),\mu^{\operatorname{as}}_{\tau}).$
Theorem 9.4
Let $G$ be almost split Kac–Moody $\mathbb{R}$-group and denote by $G(\mathbb{R})$ the rational points of $G$. Denote by $X(\mathbb{R})=G(\mathbb{R})/K(\mathbb{R})$ the symmetric space attached to $G$. Then every weak flat is contained in a standard flat. Especially
  1. 1
    standard flats are exactly the maximal flats;
  2. 2
    all weak flats are Euclidean;
  3. 3
    all weak flats are flats;
  4. 4
    $G(\mathbb{R})$ acts transitively on maximal flats.
Proof:
As in the general split case, the proof works verbatim as the proof of [FHHK20, Theorem 5.17], except claim $5$. There it is stated that the stabilizers of opposite spherical residues are reductive Lie groups. In the case of the group of rational points of an almost split Kac–Moody $\mathbb{R}$-group we get the same:
From [AB08, Proposition 6.27] it follows that the stabilizer of a spherical residue is equal to a spherical parabolic subgroup. From [Rém02, Definition 12.1.3] follows that a parabolic subgroup in $G(\mathbb{R})$ is equal to a parabolic subgroup in $G$. Using the same arguments as in the last step of the proof of 4.1, it follows that this can be interpreted as a subgroup of $\operatorname{GL}_{m}(\mathbb{R})$, $m\in\mathbb{N}$.
As in the proof of 4.10, we have to embed the involution $\Theta$ into the matrix group $\operatorname{GL}_{m}(\mathbb{R})$. According to 9.1 we know that $\Theta$ commutes with every rectified Galois automorphism $\widehat{\sigma}$ and therefore we can embed $\Theta$ on these parabolic $\mathbb{R}$-groups as transpose inverses. Thus, the arguments of claim $5$ from the proof of [FHHK20, Theorem 5.17] apply verbatim. ∎